### v1

# SpIDER: Spatially Informed Dense Embedding Retrieval for Software Issue Localization

Abstract
Retrieving code units (e.g., files, classes, functions) that are semantically relevant to a given user query, bug report, or feature request from large codebases is a fundamental challenge for LLM-based coding agents. Agentic approaches typically employ sparse retrieval methods like BM25 or dense embedding strategies to identify relevant units. While embedding-based approaches can outperform BM25 by large margins, they often lack exploration of the codebase and underutilize its underlying graph structure. To address this, we propose SpIDER (Spatially Informed Dense Embedding Retrieval), an enhanced dense retrieval approach that incorporates LLM-based reasoning over auxiliary context obtained through graph-based exploration of the codebase. Empirical results show that SpIDER consistently improves dense retrieval performance across several programming languages.
1 Introduction
Figure 1: SpIDER workflow. All functions are first ranked by semantic similarity to the issue description, and the top-$K$ functions are retrieved. From these, the top-$C$ functions (where $C\leq K$) serve as centers for spatial exploration along ‘contains’ edges. For each center, we use breath first search to explore neighboring functions within $d$ hops, considering only neighbors that also rank within the top-$N$ by semantic similarity. Then an LLM selects the most relevant neighbors for each center. Finally, selected neighbors are inserted immediately below their corresponding centers in the ranked list.
Recent work has demonstrated the potential of agents powered by Large Language Models (LLMs) to perform automated bug repair and implement basic features within large-scale software repositories (yang2024swe; aider; xia2024agentlessdemystifyingllmbasedsoftware). A critical prerequisite for effective code generation is the precise identification of relevant contextual information, specifically code units such as functions, classes, or files. This code localization task represents a foundational yet notably difficult challenge, particularly when operating at the fine-grained level of function retrieval (jimenez2024swebench; rashid2025swepolybenchmultilanguagebenchmarkrepository; zan2025multiswebenchmultilingualbenchmarkissue; chen2025locagent). Poor code retrieval directly compromises the quality of generated code, leading to incorrect or incomplete fixes, the introduction of new bugs, and substantial increases in both the computational cost and duration of the repair process. Effective code localization is thus essential for realizing the promise of autonomous software engineering agents.
Current approaches to code localization fall into two primary categories. Some recent methods exploit graph representations of code repositories alongside the reasoning capabilities of LLMs (ouyang2025repograph; chen2025locagent; jiang2025cosilsoftwareissuelocalization). These approaches primarily employ sparse retrieval strategies such as BM25 (Robertson1994OkapiATBM25) to identify relevant graph nodes or subgraphs for agent exploration. Other methods train bi-modal encoders that semantically align dense embeddings of code units with issue descriptions(fehr-etal-2025-coret; reddy2025sweranksoftwareissuelocalization). These dense retrieval methods rank code units, such as functions, classes, or files, based on their embedding similarity (typically cosine similarity) to the issue description.
However, existing approaches face three critical limitations. First, conventional contrastive learning-based retrieval methods overlook the structural context of code modules within a repository when ranking them. We observe that in software issue localization, buggy code modules are often spatially proximate within the codebase (e.g., within the same file or in neighboring functions), and top-ranked entities retrieved via dense semantic similarity frequently reside near the actual buggy modules. Consequently, relying exclusively on semantic similarity to select the top-$K$ functions may overlook relevant candidates that are spatially close to highly-ranked modules but exhibit marginally lower embedding similarity scores. Second, existing methods have been primarily evaluated on Python benchmarks (chen2025locagent; fehr-etal-2025-coret), particularly SWEBench-Lite, SWEBench-Verified, and LocBench (jimenez2024swebench; chowdhury2024introducingswebenchverified; chen2025locagent) – leaving their performance on other programming languages largely underexplored. Third, function-level retrieval, which represents the most challenging granularity compared to class-level and file-level retrieval(rashid2025swepolybenchmultilanguagebenchmarkrepository; reddy2025sweranksoftwareissuelocalization; chen2025locagent), remains largely understudied.
To address these limitations, we introduce SpIDER, a novel and simple graph-aware dense retrieval strategy for function-level code localization. SpIDER incorporates spatial locality as a complementary signal to semantic similarity, enabling the joint exploitation of both code content and repository structure within a fixed retrieval budget. Our primary focus is on retrieval methods that can augment downstream tasks, including ranking code modules, precise localization, and automated patch generation for resolving GitHub-style issues. Moreover, to facilitate comprehensive evaluation across multiple programming languages, we also introduce SpIDER-Bench for code localization, encompassing Python, Java, JavaScript, and TypeScript repositories from SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench.
Our primary contributions are as follows:
  • We propose SpIDER, a novel ane simple graph-aware dense retrieval strategy that incorporates both semantic content and code structure to determine the most relevant code functions for a given issue description under a fixed retrieval budget.
  • We introduce SpIDER-Bench, a heterogeneous graph-structured dataset for multi-language code localization. It consists of software issues from existing GitHub repositories in SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench benchmarks, covering Python, Java, JavaScript, and TypeScript, with code graphs that include source code content as node-level features.
  • We provide comprehensive empirical validation of SpIDER against existing dense retrieval methods and the popular sparse retrieval method BM25 across all four languages in SpIDER-Bench.
2 Related Work
Code Retrieval
Recent approaches like zhang2024codesage; suresh2025cornstack; feng-etal-2020-codebert; guo2021graphcodebertpretrainingcoderepresentations have proposed bi-modal encoder models to encode both code chunks as well as text embeddings in an aligned latent space. However, such models are pretrained on standard natural language to code generation tasks yielding a suboptimal performance for code retrieval based on issue descriptions. fehr-etal-2025-coret; reddy2025sweranksoftwareissuelocalization then finetune such models to the Github-style issues dataset to achieve better embedding alignment between issue descriptions and code representations.
Code Graph Construction
Given a code repository, the goal is to construct a graph that can be utilized by a downstream user (e.g., LLM agent) to understand the code structure in unison with the code content to estimate the likelihood of a module to solve a given issue based on logical reasoning and understanding of the codebase. There are various unique designs of code graphs that can be constructed from a code repository based on the definition and construction of the nodes and edges. For example, in ouyang2025repograph each node of the graph corresponds to a line of code which can of type either definition or reference. The edges between nodes capture the relation between them which can of type ‘invokes’ or ‘contains’. On the other hand, in chen2025locagent, each node corresponds to a code module (function, class, file or directory) which are connected via edges of types ‘contains’, ‘invokes’, ‘imports’ and ‘inherits’. jiang2025cosilsoftwareissuelocalization defines two types of graphs, directory-level graph (called module call graph) and file-level graph (called function call graph). Directory-level graph captures the directory structure where each node is a file which in turn can be expanded into function call graph, each node being a function. The files are connected using ‘imports’ edge whereas functions are connected using ‘invokes’ or ‘inherits’ edges.
3 Problem Definition and Notations
Given a retrieval budget $K$, issue description $Q$, and a codebase represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $n$ nodes, the task is to retrieve the top-$K$ functions that are most likely to require edits in order to resolve the issue described by $Q$. Here, $\mathcal{V}=\{v_{i}\}_{i=1}^{n}$ is the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. Each node $v_{i}\in\mathcal{V}$ represents a code entity such as a function, class, file, or directory, and entities of interest (in this case function nodes) contain the corresponding source code content. Each edge $e(v_{i},v_{j})\in\mathcal{E}$ denotes a relation between code entities, and can be one of the four edge types, ‘contains‘, ‘imports‘, ‘invokes‘ and ‘inherits‘.
Let $\mathcal{V}^{*}=\{v^{*}_{i}\}^{m}_{i=1}\subseteq\mathcal{V}$ denote the ground-truth set of $m$ relevant function nodes for issue $Q$. Let $\mathcal{V}_{K}=\{\hat{v}_{i}\}^{K}_{i=1}\subseteq\mathcal{V}$ denote the top-$K$ function nodes retrieved by a model. The objective is to maximize Recall@$K$ and Accuracy@$K$ (denoted as Acc@$K$), defined as follows:
$\text{Recall@}K=\frac{|\mathcal{V}^{*}\cap\mathcal{V}_{K}|}{|\mathcal{V}^{*}|}$
$\text{Acc@}K=\mathbb{I}[\mathcal{V}^{*}\subseteq\mathcal{V}_{K}]$
Intuitively, Recall@$K$ measures the proportion of ground truth nodes that appear in the retrieved top-$K$ list. In contrast, Acc@$K$ is a stricter metric: it is 1 for an instance only if all of its ground truth nodes are included in the top-$K$ results. At the dataset level, mean Acc@$K$ reflects the proportion of instances that achieve perfect recall (i.e., Recall@$K$ = 1).
4 Building Code Graphs
To represent a codebase as a graph, we follow chen2025locagent, and define each node as a module that can be either function, class, file or directory, connected by edges of type ‘imports’, ‘invokes’, ‘contains’ and ‘inherits’. We use python AST module [^1] https://docs.python.org/3/library/ast.html for constructing graphs for Python programming language, whereas for Java, JavaScript and TypeScript, we use tree-sitter [^2] https://github.com/tree-sitter implementation to construct similar code graphs. The graph construction starts by navigating the repository. For a directory encounter, it is added as a directory node to the graph. Similarly, if a file is encountered, it is added to the graph if it contains a class or a function, otherwise, it is ignored. Then we analyze the valid files (ending in ‘.py’ for python, ‘.java’ for Java, ‘.js’ for JavaScript and ‘.ts’ for TypeScript) further to handle various types of code lines in the file parsed by the tree-sitter or AST python module.
The core challenge in constructing graphs for Java, JavaScript and TypeScript lies in the many ways functions and classes can be defined and organized. For instance, a class might be declared in one file with no method implementations, while its methods are defined elsewhere, either in a different part of the same file or even in a completely separate file. This separation complicates the process of building a complete and accurate program graph. We explicitly handle diverse scenarios of such declaration/definitions using Tree-sitter. Additional details about the dataset and its construction can be found in section˜A.1.
Table 1: SpIDER-Bench data statistics after graph construction (GT = ground truth)
GT Type Original Dataset Python Java JavaScript TyeScript
% Insts. with Avg. GT edits % Insts. with Avg. GT edits % Insts. with Avg. GT edits % Insts. with Avg. GT edits
GT edits per inst. GT edits per inst. GT edits per inst. GT edits per inst.
Function SWE-PolyBench 85.93 3.73 70.30 5.52 53.98 1.98 29.63 2.68
Multi-SWEBench N/A N/A 59.54 2.46 81.74 3.88 27.68 2.29
SWEBench-Verified 89.00 1.82 N/A N/A N/A N/A N/A N/A
Class SWE-PolyBench 75.38 2.9 87.27 3.07 20.45 1.37 19.07 1.91
Multi-SWEBench N/A N/A 81.68 1.58 9.55 1.06 3.13 1.57
SWEBench-Verified 74.00 1.40 N/A N/A N/A N/A N/A N/A
File SWE-PolyBench 99.50 1.91 100.00 3.2 58.31 1.68 55.97 2.00
Multi-SWEBench N/A N/A 97.71 1.56 98.60 2.37 70.54 5.08
SWEBench-Verified 100.00 1.24 N/A N/A N/A N/A N/A N/A
While the graph includes four types of nodes, we focus on retrieving functions, which represent the most granular and therefore the most challenging module type. This choice is further motivated by data from the SWE-PolyBench dataset, where the majority of instances ($67\%$) contain ground truth edits at the function level, with an average of $2.78$ edits per instance rashid2025swepolybenchmultilanguagebenchmarkrepository. In contrast, only $1.42\%$ of instances involve edits at the class-only level (i.e., excluding changes to methods within the class), with an average of just $0.49$ edits per instance. This disparity highlights both the prevalence and complexity of function-level edits, highlighting the importance of accurate function-level retrieval. Additionally, since files are composed of functions and classes, effective retrieval at the function and class level naturally supports accurate file-level retrieval as well. Table 1 shows the final dataset statistics. We note down the percentage of instances retained from each dataset for all languages. The retention depends on the type of ground truth node (class, file or function) that needs to be edited to fix the issue for a given instance. If any of the ground truth nodes for an instance is a function, then the instance is retained and ignored otherwise. We also report the average edits per instance for all the languages and datasets.
5 SpIDER Framework
Algorithm 1 SpIDER
0: Code graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, issue description text $Q$, code/text dense encoder $\mathcal{F}(\cdot)$,Parameters: retrieval budget $K$, filtering threshold NN, number of centers $C$, search depth $d$
1: Compute node embeddings: $\mathcal{F}(\mathcal{V})=\{\mathcal{F}(v_{i})\}_{i=1}^{n}$
2: Compute query embedding: $\mathcal{F}(Q)$
3: Compute cosine similarities: $\mathcal{S}=\{s_{i}=f(v_{i})\cdot f(Q)\}_{i=1}^{n}$
4: Let $\mathcal{I}_{K}\subset\{1,\dots,n\}$ be the ordered set of indices of the top-$K$ nodes sorted by $s_{i}$ in descending order
5: Let $\mathcal{I}_{N}\subset\{1,\dots,n\}$ be the ordered set of indices of the top-$N$ nodes sorted by $s_{i}$ in descending order, with $N>K$
6: Let $\mathcal{V}_{K}=\{v_{i}\mid i\in\mathcal{I}_{K}\}$ and $\mathcal{V}_{N}=\{v_{i}\mid i\in\mathcal{I}_{N}\}$
7: Let $\mathcal{I}_{C}\subset\mathcal{I}_{K}$ be the set of first $C$ indices in $\mathcal{I}_{K}$
8: Let $\mathcal{V}_{C}=\{v_{i}\mid i\in\mathcal{I}_{C}\}$ {Center nodes for BFS}
9: For each node $v\in\mathcal{V}_{C}$, perform BFS in $\mathcal{G}$ up to depth $d$ and collect neighbors of type ‘function‘ that also belong to $\mathcal{V}_{N}$
10: Apply an LLM-based filtering to these neighbors and retain those most likely to resolve the issue $Q$. The LLM is provided the source code content of all the neighbors of a center node along with the issue description and asked to select only the neighbors that are most likely to solve the issue. Let $\mathcal{I}_{\text{LLM}}$ denote the set of indices selected by the LLM, consisting of the indices of the center nodes and the LLM-selected neighbors from among the BFS retrieved neighbors.
[…]
Figure 1 and Algorithm 1 provides a high-level overview of our proposed framework. Given an issue description (bug report, feature requests, etc.) and codebase, SpIDER retrieves top $K$ most relevant functions that should be modified to solve the issue. This method can be easily extended to class level as well as file level retrieval by using subset of code chunks if the entire class or file level code cannot fit into the LLM context (refer to section˜A.2 for more details). For this reason, we use the terms ‘function retrieval’ and ‘node retrieval’ interchangeably.
SpIDER’s design reflects our key observation: for instances with multiple edits, it is often the case that the edits are spatially close to each other in a code graph due to the dependencies on other functions/classes/files. Therefore, modifying a function can have cascading effects, requiring modifications to all other functions that call it, as well as to its containing class/file. Similarly, while semantically retrieved nodes offer diverse coverage of the potential candidate functions across the codebase, they may not be fully leveraging the "near miss" scenarios, where the true solution may spatially close to the top $K$ ranked functions. SpIDER tries to improve the coverage by balancing the neighborhood exploration and semantic retrieval to retrieve the most likely set of $K$ functions.
Semantic Retrieval.
The first step is to retrieve the top $K$ functions based on the semantic similarity of their dense embeddings to the issue description. We can obtain dense embeddings from encoders like CodeSAGE zhang2024codesage or an embedding model from SweRank reddy2025sweranksoftwareissuelocalization. Once function embeddings are generated in the code graph, we rank functions by cosine similarity to the issue description embedding. Next, we retrieve top $K$ functions from the semantically ranked list of functions. We call the set of node indices of the top $K$ functions ranked in the descending order of their similarity scores with the issue descriptions as $\mathcal{I}_{K}$. We also store a set of node indices of top $N$ ranked functions $\mathcal{I}_{N}$ where $N>K$. This is our primary set of retrieved functions which we aim to further improve using neighborhood exploration.
Neighborhood Exploration.
Since we have a retrieval limit of $K$, it is not affordable to store the entire neighborhood of all the nodes in $\mathcal{I}_{K}$. To strictly limit the neighbors we explore and add to the final retrieval list of the same size $K$, we restrict the neighborhood exploration to the first $C$ nodes in $\mathcal{I}_{K}$ within a $d$-hop distance from the center nodes. While we primarily focus on exploring neighbors via the ‘contains’ edge type to capture the structural organization of the codebase, SpIDER can be extended to incorporate other edge types like ‘imports’, ‘invokes’, ‘inherits’, etc. for further performance enhancements. Two functions within the same class would be 2 hops away from each other. Similarly, two functions within two separate classes in two separate files will be 4 hops away. We do not explore neighbors beyond $d$ hop distance. Increasing the depth improves coverage at the cost of increased neighbors context per center. Hence, $d$ and $C$ are hyperparameters that a user can adjust based on the downstream budget, thereby offering more flexibility and tuning capabilities in terms of retrieval performance.
For evaluation, we set $C=5$ and $N=500$ based on optimal performance observed for Python in the SWE-PolyBench dataset (see table˜4 and table˜12). These values are kept fixed across all languages and datasets to avoid hyperparameter overfitting. Additionally, we set $d=4$ to ensure that token consumption per LLM invocation remains below 10 00010\,000. While higher values of $K$ generally improve retrieval performance (see table˜13, fig.˜2), we evaluate for $K=20$ in our main results (Tables 3 and 2) to ensure fairness to the SweRank baseline, which selects $K=20$ as optimal for its downstream LLM-based reranker that improves localization at higher ranks.
Neighbor filtering
[…]
6 Experiments
Our goal is to answer the following three research hypotheses from our experiments:
H1 Can the graph structure of a codebase be leveraged to improve retrieval recall and accuracy for instances involving multiple function edits that are spatially close within the code graph?
H2 Among the top-$K$ nodes retrieved using dense embedding retrieval, do the highest-ranked nodes serve as accurate approximations of the actual ground truth locations in the code graph? If so, is it possible to reach the true nodes through local exploration of the code graph, potentially trading off some lower-ranked retrieved nodes in favor of better spatial coverage?
H3 How do existing retrieval methods, along with our proposed approach, perform across programming languages beyond Python?
To this end, we focus our study towards scenarios having more than 1 edit per instance as shown in table 1 for 4 programming languages, namely Python, Java, JavaScript and TypeScript, highlighting the robustness of the baselines and our approach.
6.1 Experimental Setup
(a)  Python
(b)  Java
(c)  JavaScript
(d)  TypeScript
Figure 2: Dense retrieval methods performance comparison on SWE-PolyBench benchmark.
Datasets
We conduct experiments on SWEBench-Verified, SWE-PolyBench and Multi-SWEBench considering patched functions as the ground truth targets for retrieval. Following suresh2025cornstack, we exclude the instances having no function level modifications in their patch. We evaluate across the Python, Java, TypeScript, and JavaScript programming languages.
Evaluation metrics
Since, the goal is to improve the coverage when retrieving $K$ nodes, we mainly use Recall@$K$ and Acc@$K$ for evaluating and comparing various methods. Here Acc@$K$ for an instance is 1 if all the ground truth targets are present in $K$ retrieved predictions. We also present Mean Reciprocal Rank (MRR) which is the inverse of the rank of the first encounter of the ground truth node in the top-$K$ retrieved list.
Hyperparameters
The hyperparameters for our algorithm are $C$, $d$, NN and $K$. We present various studies ablating on each of these parameters in the subsequent sections. Following reddy2025sweranksoftwareissuelocalization; fehr-etal-2025-coret we use $K=20$ for all the datasets to be fair to localization methods that are sensitive to the choice of this hyperparameter such as SWERankEmbed + reranker. We perform a grid search to select SpIDER’s hyperparameters on SWE-PolyBench’s Python split, resulting in $C=5$, $d=4$ and $N=500$. We keep these values constant for all our experiments. In our ablation studies, we analyze how these design choices affect the LLM’s context length requirements and invocation budget, providing deeper insights for downstream users. For all experiments, we set the LLM temperature to $0.1$.
Baselines
We compare SpIDER against state-of-the-art retrieval approaches based on dense embedding models like SweRankEmbed-Small ZS, CodeSAGE-Small ZS, reddy2025sweranksoftwareissuelocalization; zhang2024codesage and BM25 Robertson1994OkapiATBM25 as it represent a popular sparse retrieval method. Because CodeSAGE-Small ZS is not trained for Github-style problem solving tasks, we further finetune it using parameter efficient finetuning using LoRA finetuning (hu2022lora) on $2000$ instances from 3 repositories of the SWEBench training dataset. We call this baseline CodeSAGE-Small-PEFT. fehr-etal-2025-coret did not release the embedding model as well as code publicly. Hence, we could not use it as a baseline. We do not include LocAgent (chen2025locagent) due to two main drawbacks: it is extremely slow and requires significantly more number of LLM invocations as opposed to SpIDER. To give an estimate of performance of LocAgent compared to other methods, we evaluate its performance on SWEBench-Verified and on Python repositories of SWE-PolyBench benchmark table˜9. It took us more than 50 hours to complete the evaluation of LocAgent on SWEBench-Verified while SpIDER finished the evaluation in less than 2 hours. Additionally, SpIDER requires $C$ invocations of an LLM which is not deterministic for LocAgent.
We improve each baseline using spatial information from code graphs and demonstrate two retrieval approaches for each model, one using SpIDER and the other using dense retrieval with cosine similarity for dense embedding models and the standard sparse retrieval implementation for BM25.
6.2 Results
Table 2: Comparison of dense retrieval and dense retrieval + reranker/selector for $K=20$ and $N=500$, $C=5$, $d=4$. Winners per metric are highlighted in light blue.
Language Model Retrieval Dense Retrieval Only Dense + Reranker/Selector
Recall@20 Accuracy@20 MRR@20 Recall@3 Accuracy@3 MRR@3
Python SweRankEmbed-Small ZS SpIDER 0.49 0.40 0.30 0.44 0.36 0.54
DER 0.42 0.35 0.26 0.38 0.31 0.48
CodeSAGE-Small ZS SpIDER 0.27 0.21 0.13 0.23 0.18 0.31
DER 0.21 0.17 0.11 0.19 0.16 0.25
CodeSAGE-Small-PEFT SpIDER 0.37 0.27 0.24 0.32 0.25 0.42
DER 0.29 0.20 0.21 0.26 0.20 0.38
BM25 SpISR 0.33 0.26 0.19 0.30 0.23 0.38
SR 0.31 0.24 0.18 0.27 0.22 0.35
Java SweRankEmbed-Small ZS SpIDER 0.36 0.27 0.18 0.25 0.17 0.36
DER 0.31 0.24 0.18 0.24 0.16 0.33
CodeSAGE-Small ZS SpIDER 0.23 0.17 0.10 0.16 0.09 0.26
DER 0.15 0.10 0.07 0.11 0.07 0.19
CodeSAGE-Small-PEFT SpIDER 0.28 0.18 0.15 0.20 0.14 0.28
DER 0.21 0.14 0.13 0.17 0.11 0.25
BM25 SpISR 0.16 0.10 0.10 0.12 0.08 0.20
SR 0.15 0.10 0.10 0.12 0.09 0.19
JavaScript SweRankEmbed-Small ZS SpIDER 0.60 0.52 0.18 0.43 0.37 0.43
DER 0.52 0.45 0.16 0.41 0.34 0.42
CodeSAGE-Small ZS SpIDER 0.38 0.32 0.09 0.30 0.25 0.31
DER 0.26 0.21 0.07 0.22 0.17 0.24
CodeSAGE-Small-PEFT SpIDER 0.55 0.48 0.15 0.41 0.35 0.41
DER 0.45 0.38 0.14 0.36 0.30 0.39
BM25 SpISR 0.46 0.39 0.13 0.36 0.30 0.38
SR 0.36 0.29 0.11 0.31 0.25 0.34
TypeScript SweRankEmbed-Small ZS SpIDER 0.35 0.28 0.06 0.27 0.21 0.29
DER 0.29 0.21 0.18 0.25 0.19 0.30
CodeSAGE-Small ZS SpIDER 0.16 0.10 0.03 0.14 0.09 0.19
DER 0.10 0.06 0.06 0.09 0.06 0.14
CodeSAGE-Small-PEFT SpIDER 0.25 0.18 0.04 0.22 0.16 0.27
DER 0.19 0.13 0.11 0.18 0.12 0.24
BM25 SpISR 0.11 0.09 0.02 0.10 0.07 0.11
SR 0.07 0.05 0.04 0.07 0.05 0.09
Leveraging Code Graphs Improves Retrieval
We assess the effect of graph structure on code retrieval in SWE-PolyBench, comparing standard dense embeddings to our graph-aware SpIDER across Python, TypeScript, JavaScript, and Java (Fig. 2). The top-$K$ modules are retrieved by ranking code embeddings against issue embeddings via cosine similarity. SpIDER (SweRankEmbed-Small + SpIDER), which incorporates spatially informed graph structure, consistently outperforms all baselines across languages and $K$ values, highlighting the benefit of leveraging codebase graph information. Notably, the zero-shot SweRankEmbed-Small model– trained solely on $\approx$3,300 Python repositories– remains competitive across non-Python languages, surpassing both BM25 and codesage-small-v2 in zero-shot and PEFT settings, demonstrating strong cross-lingual transfer. Overall, while semantic embeddings provide a solid foundation, explicitly modeling spatial relationships via SpIDER yields substantial additional gains in retrieval accuracy.
Table 3: Dense retrieval performance across benchmarks for $K=20$, $N=500$, $C=5$, $d=4$. Winners per metric highlighted in light blue.
Language Model Retrieval Recall@20 Accuracy@20 MRR@20
Multi-SWEBench
Java SweRankEmbed-Small ZS SpIDER 0.37 0.28 0.08
DER 0.32 0.24 0.08
CodeSAGE-Small ZS SpIDER 0.21 0.19 0.05
DER 0.15 0.13 0.03
CodeSAGE-Small-PEFT SpIDER 0.31 0.25 0.10
DER 0.23 0.17 0.09
BM25 SpISR 0.23 0.17 0.07
SR 0.22 0.17 0.05
JavaScript SweRankEmbed-Small ZS SpIDER 0.39 0.30 0.18
DER 0.31 0.23 0.15
CodeSAGE-Small ZS SpIDER 0.17 0.12 0.08
DER 0.09 0.06 0.05
CodeSAGE-Small-PEFT SpIDER 0.32 0.24 0.13
DER 0.25 0.18 0.11
BM25 SpISR 0.21 0.14 0.08
SR 0.14 0.10 0.06
TypeScript SweRankEmbed-Small ZS SpIDER 0.52 0.42 0.08
DER 0.40 0.32 0.23
CodeSAGE-Small ZS SpIDER 0.33 0.26 0.04
DER 0.26 0.21 0.13
CodeSAGE-Small-PEFT SpIDER 0.30 0.26 0.05
DER 0.19 0.15 0.12
BM25 SpISR 0.32 0.27 0.04
SR 0.24 0.21 0.11
SWEBench-Verified
Python SweRankEmbed-Small ZS SpIDER 0.61 0.56 0.32
DER 0.54 0.49 0.29
CodeSAGE-Small ZS SpIDER 0.30 0.27 0.13
DER 0.21 0.19 0.10
CodeSAGE-Small-PEFT SpIDER 0.55 0.49 0.27
DER 0.43 0.38 0.22
BM25 SpISR 0.45 0.41 0.25
SR 0.38 0.34 0.22
Improvement is Consistent Across Datasets, Metrics, % Languages
Next, we augment all baseline retrieval methods—including dense models (SweRankEmbed-Small, codesage-small-v2) and BM25 sparse retrievers– with code graph structure information and measure the resulting improvements. We perform this analysis across all datasets (SWEBench-Verified, SWE-PolyBench, Multi-SWEBench) as reported in Tab. 2 and 3.
Our results reveal that our SpIDER strategy consistently enhances retrieval performance, particularly in Recall@K and Acc@K, across all programming languages and datasets. The best-performing embedding model across datasets is SweRankEmbed-Small, and applying SpIDER further increases Recall@20 and Acc@20 by at least $12.9\%$ and $14\%$, respectively.
Fig. 6 shows that SpIDER consistently outperforms baseline DER methods across varying numbers of function-level edits when using SweRankEmbed-Small. With the LLM temperature set to 0.1, the standard deviation across multiple iterations was $\leq 0.01$ for all metrics, indicating stable performance.
MRR@$K$ Gains Depend on Graph Neighborhoods
The MRR@$K$ metric benefits from SpIDER primarily when the first relevant ground-truth item within the top-$K$ results is a neighbor of a center node. In cases where the first relevant item is a non-center node, particularly in TypeScript as shown in Tab. 3SpIDER may rank neighbors of center nodes ahead, slightly reducing MRR@$K$. Since SpIDER is not explicitly designed to optimize the precise ranking order within the top-$K$ results, improvements in MRR@$K$ over baseline DER methods are modest but still observable.
Retrieval Strategy is Agnostic to Downstream Rerankers
Our retrieval strategy is agnostic to downstream rerankers or localization steps. To validate this, we evaluate an LLM selector (similar to the reranker in reddy2025sweranksoftwareissuelocalization) that retrieves the top 3 modules from a set of $K=20$ candidates. As reported in Tab. 2, improving top-$K$ coverage with SpIDER also translates to notable gains in Recall@20 and Acc@20 for top-3 localized modules.
6.3 Ablation Experiments
Table 4: Impact of the number of center nodes ($C$) on retrieval performance using SpIDER + SweRankEmbed-Small. Increasing $C$ generally improves Recall@20 and Acc@20 across languages, while MRR@20 shows modest variation.
$\boldsymbol{C}$ Python JavaScript Java TypeScript
Recall@20 Acc@20 MRR@20 Recall@20 Acc@20 MRR@20 Recall@20 Acc@20 MRR@20 Recall@20 Acc@20 MRR@20
1 0.47 0.38 0.29 0.56 0.48 0.18 0.32 0.24 0.18 0.32 0.24 0.06
3 0.49 0.39 0.30 0.59 0.51 0.18 0.35 0.26 0.19 0.31 0.24 0.06
5 0.49 0.40 0.30 0.60 0.52 0.18 0.36 0.27 0.18 0.35 0.28 0.06
7 0.49 0.39 0.30 0.60 0.52 0.18 0.36 0.27 0.19 0.37 0.30 0.06
9 0.48 0.39 0.30 0.60 0.53 0.19 0.37 0.27 0.19 0.39 0.31 0.06
Table 5: Impact of the number of neighbors per center ($d$) on retrieval performance and LLM token usage using SpIDER + SweRankEmbed-Small for $K=20$, $N=100$, $C=3$. Increasing $d$ improves Recall@20 and Accuracy@20 across languages, while LLM input/output tokens scale with $d$. Best values per metric are highlighted.
$\boldsymbol{d}$ Python JavaScript Java TypeScript
Recall@20 Acc@20 Input Output Recall@20 Acc@20 Input Output Recall@20 Acc@20 Input Output Recall@20 Acc@20 Input Output
2 0.45 0.36 2,642 32 0.54 0.46 2,032 50 0.33 0.25 2,158 22 0.31 0.23 1,474 21
4 0.46 0.37 3,591 35 0.57 0.49 3,115 25 0.33 0.25 2,083 54 0.30 0.24 1,994 24
6 0.50 0.41 11,904 48 0.62 0.53 10,065 42 0.33 0.25 3,615 67 0.36 0.28 4,334 33
8 0.55 0.44 29,658 69 0.67 0.58 15,028 48 0.34 0.27 5,594 82 0.36 0.28 5,643 37
On the Number of Top $C$ Centers
In this section, we provide ablation studies on varying the number of top $C$ centers on SWE-PolyBench for retrieval performance. Increasing $C$ results in more neighborhood explorations and addition of relevant neighbors in the list of top-$K$ retrieved modules at the cost of dropping bottom semantically ranked modules. Furthermore, each neighborhood exploration around a center requires 1 LLM call for filtering neighbors. Hence, increasing $C$ results in increased number of LLM calls. In table˜4, we ablate on $C$ and observe the corresponding performance changes. We see that the performance saturates at $C=5$ for $N=500,d=4$ at $K=20$. Hence, we stick with $C=5$ for all our experiments. A downstream user can aim to choose a $C$ that fits in their LLM invocation limits per instance.
On $K$
The choice of $K$ for retrieval depends on the budget of the downstream application. Retrieval performance is expected to improve as $K$ increases. We further validate this claim in figures 2 and 5 where we see a monotonic improvement in performance as $K$ incremented. The downstream user can choose a $K$ that is allowed by their budget while aiming for best retrieval performance.
On depth $d$
In this section, we vary the maximum depth $d$ for breadth-first search (BFS) exploration while selecting neighbors and observe the corresponding impact on the performance. Greater depth would result in more number of neighbors per center that need to be filtered using the two-stage neighborhood filtering. In table˜5, we ablate on $d$ keeping $N=100,C=3$ and $K=20$. We see that there is notable improvement in the Recall@20 and Acc@20 performance with increasing depth $d$, however, increasing $d$ also increases the token consumption to LLM while filtering the neighbors. The downstream user can choose $d$ based on their token budget per LLM invocation.
On the Primary Neighborhood Filtering Threshold NN
In table˜12, ablate on $N$ to see it’s impact on the retrieval performance. Higher $N$ results in more neighbors passing the primary filteration and hence, more neighbors for the LLMs to sift from. We see that the performance saturates around $N=300$ for $C=3,K=20,d=4$. Hence, the choice of $N$ offers the same flexibility as $d$ to the downstream user. The user can choose $N$ based on the token limit per LLM call.
7 Conclusion
We introduced SpIDER, an neurosymbolic code retrieval framework built on dense embedding models and spatial information from graph representations of a codebase. Through rigorous empirical analysis, we demonstrated the efficacy and efficiency of our method over existing retrieval methods in the the challenging setting of function-level retrieval. SpIDER’s exploration depth can be efficiently controlled by adjusting the subgraph size during retrieval, maintaining robust performance even with compact subgraph configurations for token-constrained applications. While, in this work, we pioneer retrieval evaluation for non-python programming languages such as Java, JavaScript and TypeScript, SpIDER can be readily adapted to other programming languages as it builds on the Tree-sitter library.
Appendix A Appendix
A.1 Graph construction details
JavaScript did not have a unique class data type until 2015 as it was only introduced ECMAScript 2015 also known as ES6. Hence most of the functionality of classes were implemented through functions, causing increased variations in the definition styles of a function (e.g. a nested function with the definition of container function at some other location in the same/different file).
A.2 Extension to Class-level edits
Table 6: Class-level performance on SWE-PolyBench for $K=20$, $N=500$, $C=5$, $d=2$. Winners per metric are in bold.
Language Model Retrieval Recall@20 Accuracy@20 MRR@20
Python SweRankEmbed-Small ZS SpIDER 0.65 0.54 0.36
DER 0.59 0.48 0.45
CodeSAGE-Small ZS SpIDER 0.39 0.33 0.16
DER 0.33 0.25 0.19
CodeSAGE-Small-PEFT SpIDER 0.31 0.26 0.16
DER 0.30 0.22 0.19
BM25 SpISR 0.60 0.52 0.28
SR 0.58 0.48 0.35
Java SweRankEmbed-Small ZS SpIDER 0.48 0.37 0.34
DER 0.46 0.37 0.34
CodeSAGE-Small ZS SpIDER 0.25 0.17 0.11
DER 0.23 0.17 0.10
CodeSAGE-Small-PEFT SpIDER 0.26 0.18 0.12
DER 0.28 0.21 0.14
BM25 SpISR 0.33 0.26 0.22
SR 0.33 0.26 0.25
TypeScript SweRankEmbed-Small ZS SpIDER
DER
CodeSAGE-Small ZS SpIDER
DER
CodeSAGE-Small-PEFT SpIDER
DER
BM25 SpISR
SR
In table˜6, we demonstrate the efficacy of the proposed approach at a class level. Here, we include only Python and Java here, excluding JavaScript, as most JavaScript repositories in the SWE-PolyBench dataset lack significant use of class syntax (introduced in ECMAScript 2015), resulting in very few instances with class ground truth nodes. A ground truth edit is classified as a class-level edit (and the corresponding class as a ground truth class node), if the patch introduces a modification within any part of the class declaration or definition (e.g., constructors, destructors, or methods) to address the issue description.
A.3 Additional Experiments
Figure 3: KDE (Kernel Density Estimate) plots with bootstrapped results of SWE-PolyBench benchmark for Recall@20 performance across various retrieval methods.
Figure 4: KDE (Kernel Density Estimate) plots with bootstrapped results of SWE-PolyBench benchmark for Acc@20 performance across various retrieval methods.
Table 7: Performance of SpIDER + SweRankEmbed-Small ZS encoder for different numbers of centers with $K=20$, $N=100$, $C=3$
$\boldsymbol{d}$ Python JavaScript Java TypeScript
Recall@20 Accuracy@20 MRR@20 Recall@20 Accuracy@20 MRR@20 Recall@20 Accuracy@20 MRR@20 Recall@20 Accuracy@20 MRR@20
2 0.45 0.36 0.28 0.54 0.46 0.17 0.33 0.25 0.18 0.31 0.23 0.06
4 0.46 0.37 0.30 0.57 0.49 0.18 0.33 0.25 0.18 0.30 0.24 0.06
6 0.50 0.41 0.33 0.62 0.53 0.20 0.33 0.25 0.18 0.36 0.28 0.07
8 0.55 0.44 0.38 0.67 0.58 0.22 0.34 0.27 0.20 0.36 0.27 0.07
Table 8: SpIDER + SweRankEmbed-Small ZS encoder performance with varying neighbor filtering threshold NN for $K=20$, $C=3$, $d=4$
$\boldsymbol{N}$ Python JavaScript Java TypeScript
Recall@20 Accuracy@20 MRR@20 Recall@20 Accuracy@20 MRR@20 Recall@20 Accuracy@20 MRR@20 Recall@20 Accuracy@20 MRR@20
50 0.45 0.37 0.29 0.56 0.48 0.18 0.32 0.24 0.19 0.31 0.24 0.06
100 0.46 0.37 0.30 0.57 0.49 0.18 0.33 0.25 0.18 0.30 0.24 0.06
300 0.48 0.39 0.30 0.58 0.50 0.18 0.35 0.26 0.19 0.32 0.25 0.06
500 0.49 0.39 0.30 0.59 0.51 0.18 0.35 0.26 0.19 0.31 0.24 0.06
700 0.48 0.39 0.31 0.59 0.51 0.18 0.35 0.25 0.18 0.31 0.23 0.06
900 0.48 0.39 0.31 0.58 0.50 0.18 0.34 0.25 0.19 0.31 0.24 0.06
1000 0.48 0.39 0.31 0.59 0.51 0.18 0.36 0.27 0.19 0.32 0.24 0.06
Table 9: LocAgent performance on SWEBench-Verified and the Python-only subset of SWE-PolyBench. Running LocAgent took over 48 hours on SWEBench-Verified and approximately 24 hours on SWE-PolyBench, highlighting a key limitation of this approach. In addition to the long runtime, LocAgent also requires significantly more LLM invocations per instance compared to dense embedding retrieval and BM25, resulting in higher computational costs.
Dataset Metrics K
3 5 10 20 30 40 50 60 70 80 90 100
SWEBench-Verified Recall@K 0.18 0.26 0.41 0.54 0.62 0.68 0.69 0.71 0.72 0.73 0.74 0.74
Acc@K 0.17 0.23 0.37 0.48 0.57 0.61 0.63 0.64 0.66 0.66 0.67 0.67
MRR@K 0.12 0.14 0.16 0.17 0.17 0.17 0.17 0.17 0.17 0.17 0.18 0.18
SWE-PolyBench Recall@K 0.23 0.30 0.41 0.49 0.53 0.58 0.58 0.59 0.59 0.60 0.61 0.61
Acc@K 0.17 0.23 0.31 0.37 0.41 0.46 0.46 0.46 0.46 0.47 0.48 0.49
(Python subset) MRR@K 0.21 0.23 0.25 0.25 0.25 0.26 0.26 0.26 0.26 0.26 0.26 0.26
Table 10: Comparison of LocAgent with dense retrieval methods + reranker for Python repositories in SWE-PolyBench and SWEBench-Verified
Dataset Model Recall@3 Accuracy@3 MRR@3
SWEBench-Verified LocAgent 0.18 0.17 0.12
SweRankEmbed-Small ZS+ SpIDER 0.44 0.36 0.54
+ Reranker
CodeSAGE-Small ZS + SpIDER 0.23 0.18 0.31
+ Reranker
CodeSAGE-Small-PEFT + SpIDER 0.32 0.25 0.42
+ Reranker
BM25+ SpISR 0.30 0.23 0.38
SWE-PolyBench LocAgent 0.23 0.17 0.21
SweRankEmbed-Small ZS+ SpIDER 0.61 0.56 0.63
+ Reranker
CodeSAGE-Small ZS + SpIDER 0.– 0.– 0.–
+ Reranker
CodeSAGE-Small-PEFT + SpIDER 0.52 0.46 0.56
+ Reranker
BM25+ SpISR 0.– 0.– 0.–
+ Reranker
Table 11: SWE-PolyBench retrieval performance with 95% confidence intervals ($\mu\pm$ CI) for $K=20$, $N=500$, $C=5$, $d=4$
Language Model Retrieval Recall@20 Accuracy@20
Python SweRankEmbed-Small ZS SpIDER 0.49 $\pm$ 0.06 0.40 $\pm$ 0.07
DER 0.42 $\pm$ 0.06 0.35 $\pm$ 0.07
CodeSAGE-Small ZS SpIDER 0.27 $\pm$ 0.06 0.21 $\pm$ 0.06
DER 0.21 $\pm$ 0.06 0.17 $\pm$ 0.06
CodeSAGE-Small-PEFT SpIDER 0.37 $\pm$ 0.06 0.27 $\pm$ 0.07
DER 0.29 $\pm$ 0.06 0.20 $\pm$ 0.06
BM25 SpISR 0.33 $\pm$ 0.07 0.26 $\pm$ 0.07
SR 0.31 $\pm$ 0.06 0.24 $\pm$ 0.07
Java SweRankEmbed-Small ZS SpIDER 0.36 $\pm$ 0.08 0.27 $\pm$ 0.08
DER 0.31 $\pm$ 0.07 0.24 $\pm$ 0.08
CodeSAGE-Small ZS SpIDER 0.23 $\pm$ 0.07 0.17 $\pm$ 0.07
DER 0.15 $\pm$ 0.06 0.10 $\pm$ 0.05
CodeSAGE-Small-PEFT SpIDER 0.28 $\pm$ 0.08 0.18 $\pm$ 0.07
DER 0.21 $\pm$ 0.07 0.14 $\pm$ 0.06
BM25 SpISR 0.16 $\pm$ 0.06 0.10 $\pm$ 0.05
SR 0.15 $\pm$ 0.06 0.10 $\pm$ 0.05
JavaScript SweRankEmbed-Small ZS SpIDER 0.60 $\pm$ 0.04 0.52 $\pm$ 0.04
DER 0.52 $\pm$ 0.04 0.45 $\pm$ 0.04
CodeSAGE-Small ZS SpIDER 0.38 $\pm$ 0.04 0.32 $\pm$ 0.04
DER 0.26 $\pm$ 0.04 0.21 $\pm$ 0.04
CodeSAGE-Small-PEFT SpIDER 0.55 $\pm$ 0.04 0.48 $\pm$ 0.05
DER 0.45 $\pm$ 0.04 0.38 $\pm$ 0.04
BM25 SpISR 0.46 $\pm$ 0.04 0.39 $\pm$ 0.04
SR 0.36 $\pm$ 0.04 0.29 $\pm$ 0.04
TypeScript SweRankEmbed-Small ZS SpIDER 0.35 $\pm$ 0.06 0.28 $\pm$ 0.05
DER 0.29 $\pm$ 0.05 0.21 $\pm$ 0.05
CodeSAGE-Small ZS SpIDER 0.16 $\pm$ 0.04 0.10 $\pm$ 0.04
DER 0.10 $\pm$ 0.04 0.06 $\pm$ 0.04
CodeSAGE-Small-PEFT SpIDER 0.25 $\pm$ 0.05 0.18 $\pm$ 0.05
DER 0.19 $\pm$ 0.05 0.13 $\pm$ 0.05
BM25 SpISR 0.11 $\pm$ 0.04 0.09 $\pm$ 0.04
SR 0.07 $\pm$ 0.03 0.05 $\pm$ 0.03
Table 12: SpIDER + SweRankEmbed-Small performance with varying neighbor filtering threshold NN for $K=20$, $C=3$, $d=4$. Increasing NN improves coverage and Recall@20 with modest increases in LLM token usage. Avg. tokens indicate average consumption per LLM call; number of calls equals $C$. Best Recall and Accuracy per language are highlighted.
$\boldsymbol{N}$ Python JavaScript Java TypeScript
Recall Accuracy Tokens Recall Accuracy Tokens Recall Accuracy Tokens Recall Accuracy Tokens
(Input / Output) (Input / Output) (Input / Output) (Input / Output)
50 0.45 0.37 3108 / 33 0.56 0.48 2383 / 23 0.32 0.24 1845 / 49 0.31 0.24 1763 / 23
100 0.46 0.37 3591 / 35 0.57 0.49 3115 / 25 0.33 0.25 2083 / 54 0.30 0.24 1994 / 24
300 0.48 0.39 4848 / 37 0.58 0.50 5119 / 29 0.35 0.26 2644 / 56 0.32 0.25 2459 / 26
500 0.49 0.39 5708 / 39 0.59 0.51 6775 / 30 0.35 0.26 2978 / 57 0.31 0.24 2687 / 26
700 0.48 0.39 6361 / 39 0.59 0.51 7998 / 31 0.35 0.25 3203 / 59 0.31 0.23 2843 / 27
900 0.48 0.39 6903 / 40 0.58 0.50 8780 / 32 0.35 0.26 3343 / 58 0.31 0.24 2943 / 27
1000 0.48 0.39 7180 / 40 0.59 0.51 9098 / 32 0.36 0.27 3441 / 59 0.32 0.25 3001 / 27
Table 13: Effect of retrieval budget $K$ on SpIDER + SweRankEmbed-Small performance for $N=500$, $C=5$, $d=4$. Increasing $K$ consistently improves Recall@K and Accuracy@K, while MRR@K remains stable. Best values per metric are highlighted.
K Python Java JavaScript TypeScript
Recall Accuracy MRR Recall Accuracy MRR Recall Accuracy MRR Recall Accuracy MRR
3 0.31 0.25 0.27 0.17 0.09 0.17 0.32 0.26 0.16 0.17 0.12 0.06
5 0.39 0.32 0.30 0.24 0.16 0.18 0.39 0.34 0.17 0.19 0.14 0.06
10 0.44 0.36 0.30 0.33 0.23 0.19 0.49 0.41 0.17 0.29 0.23 0.06
20 0.47 0.39 0.29 0.35 0.26 0.19 0.59 0.51 0.18 0.35 0.28 0.06
30 0.52 0.42 0.30 0.37 0.27 0.19 0.62 0.54 0.18 0.40 0.31 0.06
40 0.52 0.43 0.30 0.38 0.28 0.19 0.65 0.57 0.18 0.42 0.33 0.07
50 0.54 0.44 0.30 0.42 0.31 0.19 0.68 0.59 0.18 0.42 0.33 0.06
60 0.58 0.47 0.31 0.43 0.32 0.19 0.70 0.61 0.18 0.45 0.36 0.06
70 0.58 0.46 0.30 0.45 0.34 0.19 0.72 0.63 0.18 0.47 0.38 0.07
80 0.60 0.48 0.30 0.46 0.34 0.19 0.73 0.64 0.18 0.48 0.38 0.07
90 0.61 0.50 0.30 0.46 0.34 0.19 0.75 0.67 0.18 0.48 0.39 0.06
100 0.62 0.50 0.30 0.46 0.34 0.19 0.77 0.68 0.18 0.52 0.43 0.07
Table 14: SweRankEmbed-Small zero-shot performance across varying retrieval budgets $K$ for $N=500$, $C=5$, $d=4$. Best Recall and Accuracy per language are highlighted.
K Python Java JavaScript TypeScript
Recall Accuracy MRR Recall Accuracy MRR Recall Accuracy MRR Recall Accuracy MRR
3 0.24 0.19 0.27 0.18 0.12 0.24 0.26 0.22 0.27 0.11 0.07 0.15
5 0.31 0.25 0.29 0.26 0.19 0.26 0.33 0.27 0.28 0.16 0.12 0.16
10 0.36 0.30 0.30 0.30 0.22 0.26 0.43 0.37 0.30 0.20 0.14 0.17
20 0.42 0.35 0.30 0.31 0.24 0.26 0.52 0.45 0.30 0.29 0.21 0.18
30 0.47 0.39 0.30 0.33 0.25 0.26 0.57 0.49 0.31 0.34 0.26 0.18
40 0.49 0.40 0.30 0.35 0.26 0.27 0.61 0.53 0.31 0.37 0.29 0.18
50 0.51 0.42 0.31 0.39 0.29 0.27 0.65 0.56 0.31 0.38 0.30 0.18
60 0.55 0.45 0.31 0.41 0.30 0.27 0.67 0.58 0.31 0.41 0.32 0.18
70 0.56 0.45 0.31 0.43 0.32 0.27 0.68 0.59 0.31 0.42 0.33 0.18
80 0.57 0.46 0.31 0.43 0.32 0.27 0.70 0.61 0.31 0.43 0.34 0.18
90 0.58 0.47 0.31 0.43 0.32 0.27 0.72 0.63 0.31 0.46 0.36 0.18
100 0.59 0.48 0.31 0.44 0.33 0.27 0.73 0.64 0.31 0.48 0.39 0.18
Figure 5: SpIDER vs DER performance at various values of $K$ on SWE-PolyBench benchmark for SweRankEmbed-Small ZS embedding model, $N=500$, $C=5$, $d=4$.
Figure 6: Comparing retrieval Recall@20 and Accuracy vs. number of edits for DER vs SpIDER on SweRankEmbed-Small ZS using SWE-PolyBench across various languages for $N=500$, $C=5$, $d=4$. Top row: Recall@20. Bottom row: Accuracy.

### v2

@@
 Abstract
-Retrieving code units (e.g., files, classes, functions) that are semantically relevant to a given user query, bug report, or feature request from large codebases is a fundamental challenge for LLM-based coding agents.
-Agentic approaches typically employ sparse retrieval methods like BM25 or dense embedding strategies to identify relevant units.
-While embedding-based approaches can outperform BM25 by large margins, they often lack exploration of the codebase and underutilize its underlying graph structure.
-To address this, we propose SpIDER (Spatially Informed Dense Embedding Retrieval), an enhanced dense retrieval approach that incorporates LLM-based reasoning over auxiliary context obtained through graph-based exploration of the codebase.
-Empirical results show that SpIDER consistently improves dense retrieval performance across several programming languages.
+Retrieving code functions, classes or files that are relevant in order to solve a given user query, bug report or feature request from large codebases is a fundamental challenge for Large Language Model (LLM)-based coding agents.
+Agentic approaches typically employ sparse retrieval methods like BM25 or dense embedding strategies to identify semantically relevant units.
+While embedding-based approaches can outperform BM25 by large margins, they often don’t take into consideration the underlying graph-structured characteristics of the codebase.
+To address this, we propose SpIDER (Spatially Informed Dense Embedding Retrieval), an enhanced dense retrieval approach that integrates LLM-based reasoning along with auxiliary information obtained from graph-based exploration of the codebase.
+We further introduce SpIDER-Bench, a graph-structured evaluation benchmark curated from SWE-PolyBench, SWEBench-Verified and Multi-SWEBench, spanning codebases from Python, Java, JavaScript and TypeScript programming languages.
+Empirical results show that SpIDER consistently improves dense retrieval performance by at least $13\%$ across programming languages and benchmarks in SpIDER-Bench .
 1 Introduction
 Figure 1: SpIDER workflow.
-All functions are first ranked by semantic similarity to the issue description, and the top-$K$ functions are retrieved.
-From these, the top-$C$ functions (where $C\leq K$) serve as centers for spatial exploration along ‘contains’ edges.
-For each center, we use breath first search to explore neighboring functions within $d$ hops, considering only neighbors that also rank within the top-$N$ by semantic similarity.
-Then an LLM selects the most relevant neighbors for each center.
-Finally, selected neighbors are inserted immediately below their corresponding centers in the ranked list.
-Recent work has demonstrated the potential of agents powered by Large Language Models (LLMs) to perform automated bug repair and implement basic features within large-scale software repositories (yang2024swe; aider; xia2024agentlessdemystifyingllmbasedsoftware).
+SpIDER augments standard dense embedding retrieval with graph-aware spatial exploration and LLM-based reasoning to improve function-level code localization under a fixed retrieval (K).
+Given an issue description, all functions are first ranked by semantic similarity, and the top-K candidates are retrieved.
+The top-C (C<K) functions serve as centers (func 1) for structured neighborhood exploration along contains edges in the code graph, where neighboring functions within d hops are considered if they also rank within the top-N semantically (funcs 2,3,4).
+An LLM then filters these candidates to identify functions that are likely relevant but under-ranked by semantic similarity alone (func3).
+The selected neighbors are inserted immediately below their corresponding centers while discarding an equal number of bottom ranked functions in the initial top-K list, thereby increasing coverage of structurally related buggy functions without increasing K.
+Further algorithmic details are provided in section˜A.7 (algorithm˜1).
+Recent work has demonstrated the potential of agents powered by Large Language Models (LLMs) to perform automated bug repair and implement basic features within large-scale software repositories (Yang et al., 2024; Gauthier, 2024; Xia et al., 2024).
 A critical prerequisite for effective code generation is the precise identification of relevant contextual information, specifically code units such as functions, classes, or files.
-This code localization task represents a foundational yet notably difficult challenge, particularly when operating at the fine-grained level of function retrieval (jimenez2024swebench; rashid2025swepolybenchmultilanguagebenchmarkrepository; zan2025multiswebenchmultilingualbenchmarkissue; chen2025locagent).
-Poor code retrieval directly compromises the quality of generated code, leading to incorrect or incomplete fixes, the introduction of new bugs, and substantial increases in both the computational cost and duration of the repair process.
-Effective code localization is thus essential for realizing the promise of autonomous software engineering agents.
+This code localization task represents a foundational yet notably difficult challenge, particularly when operating at the fine-grained level of function retrieval (Jimenez et al., 2024; Rashid et al., 2025; Zan et al., 2025; Chen et al., 2025).
+Poor code retrieval directly compromises the quality of generated code, leading to incorrect or incomplete fixes, the introduction of new bugs, and substantial increase in both the computational cost and duration of repair.
+Thus, effective code localization is essential to realize the promise of autonomous software engineering agents.
 Current approaches to code localization fall into two primary categories.
-Some recent methods exploit graph representations of code repositories alongside the reasoning capabilities of LLMs (ouyang2025repograph; chen2025locagent; jiang2025cosilsoftwareissuelocalization).
-These approaches primarily employ sparse retrieval strategies such as BM25 (Robertson1994OkapiATBM25) to identify relevant graph nodes or subgraphs for agent exploration.
-Other methods train bi-modal encoders that semantically align dense embeddings of code units with issue descriptions(fehr-etal-2025-coret; reddy2025sweranksoftwareissuelocalization).
-These dense retrieval methods rank code units, such as functions, classes, or files, based on their embedding similarity (typically cosine similarity) to the issue description.
-However, existing approaches face three critical limitations.
-First, conventional contrastive learning-based retrieval methods overlook the structural context of code modules within a repository when ranking them.
+Some methods exploit graph representations of code repositories alongside the reasoning capabilities of LLMs (Ouyang et al., 2025; Chen et al., 2025; Jiang et al., 2025).
+These approaches mainly employ sparse retrieval strategies such as BM25 (Robertson et al., 1994) to identify relevant graph nodes or subgraphs for agent exploration.
+Other methods train bimodal encoders using contrastive objective to semantically align dense embeddings of code units with issue descriptions (Fehr et al., 2025; Reddy et al., 2025).
+These dense retrieval methods rank code units based on their embedding similarity to the issue description.
+These existing approaches face three critical limitations.
+First, conventional contrastive learning-based retrieval methods overlook the structural context of code modules within a repository when ranking them (Reddy et al., 2025; Zhang et al., 2024; Suresh et al., 2025).
 We observe that in software issue localization, buggy code modules are often spatially proximate within the codebase (e.g., within the same file or in neighboring functions), and top-ranked entities retrieved via dense semantic similarity frequently reside near the actual buggy modules.
 Consequently, relying exclusively on semantic similarity to select the top-$K$ functions may overlook relevant candidates that are spatially close to highly-ranked modules but exhibit marginally lower embedding similarity scores.
-Second, existing methods have been primarily evaluated on Python benchmarks (chen2025locagent; fehr-etal-2025-coret), particularly SWEBench-Lite, SWEBench-Verified, and LocBench (jimenez2024swebench; chowdhury2024introducingswebenchverified; chen2025locagent) – leaving their performance on other programming languages largely underexplored.
-Third, function-level retrieval, which represents the most challenging granularity compared to class-level and file-level retrieval(rashid2025swepolybenchmultilanguagebenchmarkrepository; reddy2025sweranksoftwareissuelocalization; chen2025locagent), remains largely understudied.
-To address these limitations, we introduce SpIDER, a novel and simple graph-aware dense retrieval strategy for function-level code localization.
+Second, existing methods have been primarily evaluated on Python (Chen et al., 2025; Fehr et al., 2025), particularly SWEBench-Lite, SWEBench-Verified, and LocBench (Jimenez et al., 2024; Chowdhury et al., 2024b; Chen et al., 2025) – leaving their performance on other programming languages largely underexplored.
+Third, function-level retrieval, which represents the most challenging granularity compared to class-level and file-level retrieval(Rashid et al., 2025; Reddy et al., 2025; Chen et al., 2025), remains largely understudied.
+To address these limitations, we introduce SpIDER, a simple graph-aware dense retrieval strategy for function-level code localization.
 SpIDER incorporates spatial locality as a complementary signal to semantic similarity, enabling the joint exploitation of both code content and repository structure within a fixed retrieval budget.
-Our primary focus is on retrieval methods that can augment downstream tasks, including ranking code modules, precise localization, and automated patch generation for resolving GitHub-style issues.
-Moreover, to facilitate comprehensive evaluation across multiple programming languages, we also introduce SpIDER-Bench for code localization, encompassing Python, Java, JavaScript, and TypeScript repositories from SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench.
-Our primary contributions are as follows:
-  • We propose SpIDER, a novel ane simple graph-aware dense retrieval strategy that incorporates both semantic content and code structure to determine the most relevant code functions for a given issue description under a fixed retrieval budget.
-  • We introduce SpIDER-Bench, a heterogeneous graph-structured dataset for multi-language code localization.
-It consists of software issues from existing GitHub repositories in SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench benchmarks, covering Python, Java, JavaScript, and TypeScript, with code graphs that include source code content as node-level features.
+Our primary focus is on retrieval methods that can augment downstream tasks, including ranking code modules, precise localization, and automated patch generation to resolve GitHub-style issues.
+To facilitate comprehensive evaluation across multiple programming languages, we also introduce SpIDER-Bench for code localization, encompassing Python, Java, JavaScript, and TypeScript repositories from SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench.
+Our primary contributions are as follows: (1) We propose SpIDER, a novel and simple graph-aware dense retrieval strategy that incorporates both semantic content and code structure to determine the most relevant functions for any given issue description under a fixed retrieval budget.
+(2) We introduce SpIDER-Bench, a heterogeneous graph-structured benchmark for multi-language code localization.
+It comprises of software issues from existing GitHub repositories in SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench, covering Python, Java, JavaScript, and TypeScript, with code graphs that include source code content as node-level features.
 We provide comprehensive empirical validation of SpIDER against existing dense retrieval methods and the popular sparse retrieval method BM25 across all four languages in SpIDER-Bench.
 2 Related Work
-Code Retrieval
-Recent approaches like zhang2024codesage; suresh2025cornstack; feng-etal-2020-codebert; guo2021graphcodebertpretrainingcoderepresentations have proposed bi-modal encoder models to encode both code chunks as well as text embeddings in an aligned latent space.
-However, such models are pretrained on standard natural language to code generation tasks yielding a suboptimal performance for code retrieval based on issue descriptions.
-fehr-etal-2025-coret; reddy2025sweranksoftwareissuelocalization then finetune such models to the Github-style issues dataset to achieve better embedding alignment between issue descriptions and code representations.
-Code Graph Construction
-Given a code repository, the goal is to construct a graph that can be utilized by a downstream user (e.g., LLM agent) to understand the code structure in unison with the code content to estimate the likelihood of a module to solve a given issue based on logical reasoning and understanding of the codebase.
-There are various unique designs of code graphs that can be constructed from a code repository based on the definition and construction of the nodes and edges.
-For example, in ouyang2025repograph each node of the graph corresponds to a line of code which can of type either definition or reference.
-The edges between nodes capture the relation between them which can of type ‘invokes’ or ‘contains’.
-On the other hand, in chen2025locagent, each node corresponds to a code module (function, class, file or directory) which are connected via edges of types ‘contains’, ‘invokes’, ‘imports’ and ‘inherits’.
-jiang2025cosilsoftwareissuelocalization defines two types of graphs, directory-level graph (called module call graph) and file-level graph (called function call graph).
-Directory-level graph captures the directory structure where each node is a file which in turn can be expanded into function call graph, each node being a function.
-The files are connected using ‘imports’ edge whereas functions are connected using ‘invokes’ or ‘inherits’ edges.
+Code Retrieval and Software Issue Localization  Code retrieval refers to the problem of identifying relevant code locations from a codebase that are responsible for a software issue.
+Classical approaches are rooted in information retrieval, leveraging lexical or semantic similarity to produce a ranked list of candidate code snippets.
+Consequently, many existing retrieval systems, including some recent ones, rely on sparse retrievers such as BM25 (Robertson et al., 1994).
+While BM25 indexing is computationally cheaper than dense embedding generation and vector similarity search, it typically underperforms in retrieval quality when rich semantic representations are available (Reddy et al., 2025; Fehr et al., 2025).
+Recent dense retrieval approaches propose bimodal encoder models to encode both code chunks as well as issue embeddings in an aligned latent space (Zhang et al., 2024; Suresh et al., 2025; Feng et al., 2020; Guo et al., 2021).
+However, these models are typically pretrained on generic natural language–to–code objectives, which leads to suboptimal performance for issue-driven code retrieval.
+To address this gap, Fehr et al.
+(2025) and Reddy et al.
+(2025) introduce bimodal encoders fine-tuned on GitHub-style issue resolution datasets such as SWEBench (Jimenez et al., 2024), resulting in improved alignment between issue and code representations.
+Despite these gains, SWEBench is limited to Python repositories, and no comparably large multilingual dataset currently exists to support training such encoders across diverse programming languages.
+Consequently, as real-world codebases span multiple languages, these fine-tuned dense retrievers exhibit degraded performance under domain shifts, reducing their robustness and reliability.
+We argue that incorporating the commonsense reasoning capabilities of large language models (LLMs), together with structural signals derived from code graphs, introduces more invariant representations that are less sensitive to surface-level language artifacts.
+Augmenting dense retrievers with these components helps mitigate spurious correlations and improves generalization.
+We empirically validate this hypothesis on multilingual benchmarks, including SWE-PolyBench (Rashid et al., 2025), Multi-SWEBench (Zan et al., 2025), and SWEBench-Verified (Chowdhury et al., 2024a).
+LLM-based retrieval methods  Recent work has shown that large language models (LLMs) are highly effective at localizing and resolving GitHub-style software issues, largely due to their strong reasoning capabilities (Kang et al., 2024; Xia et al., 2024; Luo et al., 2024; Örwall, 2024).
+In the context of code retrieval, Xia et al.
+(2024) introduce a simple hierarchical localization strategy driven directly by an LLM, whereas Ouyang et al.
+(2025) employ an intermediate sparse BM25 retrieval stage to identify relevant files prior to code generation.
+Similarly, Chen et al.
+(2025) leverage inverted BM25 indexing to match issue-description keywords with node identifiers or code chunks.
+Both Ouyang et al.
+(2025) and Chen et al.
+(2025) further enable LLM-guided traversal of local subgraphs around the retrieved files or nodes, using edges defined by their respective code graph representations.
+In contrast, Reddy et al.
+(2025) focus on re-ranking densely retrieved code chunks using LLMs, with the goal of enhancing semantic retrieval through reasoning.
+Although this approach improves retrieval performance, it does not exploit the structural signals encoded in code graphs.
+We aim to bridge this gap by jointly leveraging LLM reasoning, semantic embeddings, and the additional contextual information captured by code graphs, thereby improving overall retrieval effectiveness.
+Code Graph Construction  Several recent approaches leverage explicit code graph representations to expose structural information that is otherwise difficult to capture with purely textual or embedding-based methods.
+The specific utility of such graphs depends critically on how nodes and relations are defined.
+For example, Ouyang et al.
+(2025) construct graphs at the line-of-code level, where nodes correspond to variable or module definitions and references, and edges encode invokes and contains relationships.
+While this design enables fine-grained structural reasoning, it scales poorly with repository size, leading to dense and complex graphs that are challenging for LLM-based agents to traverse at inference time.
+To improve scalability, other works adopt coarser graph abstractions.
+Jiang et al.
+(2025) propose two complementary graph types: a module call graph, connecting files via imports, and a function call graph, linking functions and classes through invokes and inherits relations.
+More recently, Chen et al.
+(2025) introduce a unified graph formulation in which nodes represent functions, classes, files, or directories, and edges capture contains’, invokes’, imports’, and ‘inherits’ relationships.
+This design strikes a balance between expressivity and tractability, enabling structural reasoning at multiple hierarchical levels while remaining amenable to LLM-guided traversal.
+Following this line of work, we adopt their graph construction, as it naturally supports retrieval at the file, class, and function levels—granularities that are common across programming languages and well aligned with multilingual code retrieval settings.
 3 Problem Definition and Notations
-Given a retrieval budget $K$, issue description $Q$, and a codebase represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $n$ nodes, the task is to retrieve the top-$K$ functions that are most likely to require edits in order to resolve the issue described by $Q$.
-Here, $\mathcal{V}=\{v_{i}\}_{i=1}^{n}$ is the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges.
-Each node $v_{i}\in\mathcal{V}$ represents a code entity such as a function, class, file, or directory, and entities of interest (in this case function nodes) contain the corresponding source code content.
-Each edge $e(v_{i},v_{j})\in\mathcal{E}$ denotes a relation between code entities, and can be one of the four edge types, ‘contains‘, ‘imports‘, ‘invokes‘ and ‘inherits‘.
-Let $\mathcal{V}^{*}=\{v^{*}_{i}\}^{m}_{i=1}\subseteq\mathcal{V}$ denote the ground-truth set of $m$ relevant function nodes for issue $Q$.
-Let $\mathcal{V}_{K}=\{\hat{v}_{i}\}^{K}_{i=1}\subseteq\mathcal{V}$ denote the top-$K$ function nodes retrieved by a model.
-The objective is to maximize Recall@$K$ and Accuracy@$K$ (denoted as Acc@$K$), defined as follows:
-$\text{Recall@}K=\frac{|\mathcal{V}^{*}\cap\mathcal{V}_{K}|}{|\mathcal{V}^{*}|}$
-$\text{Acc@}K=\mathbb{I}[\mathcal{V}^{*}\subseteq\mathcal{V}_{K}]$
-Intuitively, Recall@$K$ measures the proportion of ground truth nodes that appear in the retrieved top-$K$ list.
-In contrast, Acc@$K$ is a stricter metric: it is 1 for an instance only if all of its ground truth nodes are included in the top-$K$ results.
-At the dataset level, mean Acc@$K$ reflects the proportion of instances that achieve perfect recall (i.e., Recall@$K$ = 1).
-4 Building Code Graphs
-To represent a codebase as a graph, we follow chen2025locagent, and define each node as a module that can be either function, class, file or directory, connected by edges of type ‘imports’, ‘invokes’, ‘contains’ and ‘inherits’.
-We use python AST module [^1] https://docs.python.org/3/library/ast.html for constructing graphs for Python programming language, whereas for Java, JavaScript and TypeScript, we use tree-sitter [^2] https://github.com/tree-sitter implementation to construct similar code graphs.
-The graph construction starts by navigating the repository.
-For a directory encounter, it is added as a directory node to the graph.
-Similarly, if a file is encountered, it is added to the graph if it contains a class or a function, otherwise, it is ignored.
-Then we analyze the valid files (ending in ‘.py’ for python, ‘.java’ for Java, ‘.js’ for JavaScript and ‘.ts’ for TypeScript) further to handle various types of code lines in the file parsed by the tree-sitter or AST python module.
-The core challenge in constructing graphs for Java, JavaScript and TypeScript lies in the many ways functions and classes can be defined and organized.
-For instance, a class might be declared in one file with no method implementations, while its methods are defined elsewhere, either in a different part of the same file or even in a completely separate file.
-This separation complicates the process of building a complete and accurate program graph.
-We explicitly handle diverse scenarios of such declaration/definitions using Tree-sitter.
-Additional details about the dataset and its construction can be found in section˜A.1.
-Table 1: SpIDER-Bench data statistics after graph construction (GT = ground truth)
+We formalize the function-level code retrieval task and its evaluation, applicable to any retrieval method operating over structured code representations.
+Given a codebase and an issue description $Q$, the goal is to retrieve the top-$K$ functions that are most likely to require edits to resolve the issue.
+We represent the codebase as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $n$ nodes, where $\mathcal{V}={v_{i}}_{i=1}^{n}$ denotes the set of code entities and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ encodes structural relations.
+Each node $v\in\mathcal{V}$ corresponds to a code entity, such as a function, class, file, or directory; and edges represent relationships of types contains, invokes, imports, and inherits.
+Let $\mathcal{V}^{*}={v_{i}^{*}}_{i=1}^{m}\subseteq\mathcal{V}$ denote the ground-truth set of $m$ relevant function nodes for issue $Q$.
+Scoring Function  A bi-modal encoder $\mathcal{F}(\cdot)$ maps both code units and issue descriptions to a shared embedding space. The relevance score of node $v$ for issue $Q$ is:
+$s_{Q}(v)=\cos\bigl(\mathcal{F}(v),\mathcal{F}(Q)\bigr)$
+Baseline Dense Retrieval  Standard dense embedding retrieval (DER) ranks all nodes by their relevance scores and returns the top-$K$ candidates:
+$\mathcal{S}_{K}(Q):=\underset{v\in\mathcal{V}}{\mathrm{arg\,top}_{K}}\;s_{Q}(v)$
+where $\mathrm{arg\,top}_{K}$ returns $K$ nodes with top-$K$ scores $s_{Q}(\cdot)$. We use notation $\mathcal{S}_{K}$ and $\mathcal{S}_{K}(Q)$ interchangeably.
+4 Methodology
+Table 1: SpIDER-Bench data statistics after graph construction (GT = ground truth).
+It summarizes the distribution and granularity of ground-truth edits across languages and datasets, showing that function-level edits dominate across benchmarks while exhibiting substantial variation in edit frequency and localization difficulty across programming languages.
 GT Type Original Dataset Python Java JavaScript TyeScript
@@
 SWEBench-Verified 100.00 1.24 N/A N/A N/A N/A N/A N/A
-While the graph includes four types of nodes, we focus on retrieving functions, which represent the most granular and therefore the most challenging module type.
-This choice is further motivated by data from the SWE-PolyBench dataset, where the majority of instances ($67\%$) contain ground truth edits at the function level, with an average of $2.78$ edits per instance rashid2025swepolybenchmultilanguagebenchmarkrepository.
-In contrast, only $1.42\%$ of instances involve edits at the class-only level (i.e., excluding changes to methods within the class), with an average of just $0.49$ edits per instance.
-This disparity highlights both the prevalence and complexity of function-level edits, highlighting the importance of accurate function-level retrieval.
-Additionally, since files are composed of functions and classes, effective retrieval at the function and class level naturally supports accurate file-level retrieval as well.
-Table 1 shows the final dataset statistics.
-We note down the percentage of instances retained from each dataset for all languages.
-The retention depends on the type of ground truth node (class, file or function) that needs to be edited to fix the issue for a given instance.
-If any of the ground truth nodes for an instance is a function, then the instance is retained and ignored otherwise.
-We also report the average edits per instance for all the languages and datasets.
-5 SpIDER Framework
-Algorithm 1 SpIDER
-0: Code graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, issue description text $Q$, code/text dense encoder $\mathcal{F}(\cdot)$,Parameters: retrieval budget $K$, filtering threshold NN, number of centers $C$, search depth $d$
-1: Compute node embeddings: $\mathcal{F}(\mathcal{V})=\{\mathcal{F}(v_{i})\}_{i=1}^{n}$
-2: Compute query embedding: $\mathcal{F}(Q)$
-3: Compute cosine similarities: $\mathcal{S}=\{s_{i}=f(v_{i})\cdot f(Q)\}_{i=1}^{n}$
-4: Let $\mathcal{I}_{K}\subset\{1,\dots,n\}$ be the ordered set of indices of the top-$K$ nodes sorted by $s_{i}$ in descending order
-5: Let $\mathcal{I}_{N}\subset\{1,\dots,n\}$ be the ordered set of indices of the top-$N$ nodes sorted by $s_{i}$ in descending order, with $N>K$
-6: Let $\mathcal{V}_{K}=\{v_{i}\mid i\in\mathcal{I}_{K}\}$ and $\mathcal{V}_{N}=\{v_{i}\mid i\in\mathcal{I}_{N}\}$
-7: Let $\mathcal{I}_{C}\subset\mathcal{I}_{K}$ be the set of first $C$ indices in $\mathcal{I}_{K}$
-8: Let $\mathcal{V}_{C}=\{v_{i}\mid i\in\mathcal{I}_{C}\}$ {Center nodes for BFS}
-9: For each node $v\in\mathcal{V}_{C}$, perform BFS in $\mathcal{G}$ up to depth $d$ and collect neighbors of type ‘function‘ that also belong to $\mathcal{V}_{N}$
-10: Apply an LLM-based filtering to these neighbors and retain those most likely to resolve the issue $Q$.
-The LLM is provided the source code content of all the neighbors of a center node along with the issue description and asked to select only the neighbors that are most likely to solve the issue.
-Let $\mathcal{I}_{\text{LLM}}$ denote the set of indices selected by the LLM, consisting of the indices of the center nodes and the LLM-selected neighbors from among the BFS retrieved neighbors.
-[…]
-Figure 1 and Algorithm 1 provides a high-level overview of our proposed framework.
-Given an issue description (bug report, feature requests, etc.) and codebase, SpIDER retrieves top $K$ most relevant functions that should be modified to solve the issue.
-This method can be easily extended to class level as well as file level retrieval by using subset of code chunks if the entire class or file level code cannot fit into the LLM context (refer to section˜A.2 for more details).
-For this reason, we use the terms ‘function retrieval’ and ‘node retrieval’ interchangeably.
-SpIDER’s design reflects our key observation: for instances with multiple edits, it is often the case that the edits are spatially close to each other in a code graph due to the dependencies on other functions/classes/files.
-Therefore, modifying a function can have cascading effects, requiring modifications to all other functions that call it, as well as to its containing class/file.
-Similarly, while semantically retrieved nodes offer diverse coverage of the potential candidate functions across the codebase, they may not be fully leveraging the "near miss" scenarios, where the true solution may spatially close to the top $K$ ranked functions.
-SpIDER tries to improve the coverage by balancing the neighborhood exploration and semantic retrieval to retrieve the most likely set of $K$ functions.
-Semantic Retrieval.
-The first step is to retrieve the top $K$ functions based on the semantic similarity of their dense embeddings to the issue description.
-We can obtain dense embeddings from encoders like CodeSAGE zhang2024codesage or an embedding model from SweRank reddy2025sweranksoftwareissuelocalization.
-Once function embeddings are generated in the code graph, we rank functions by cosine similarity to the issue description embedding.
-Next, we retrieve top $K$ functions from the semantically ranked list of functions.
-We call the set of node indices of the top $K$ functions ranked in the descending order of their similarity scores with the issue descriptions as $\mathcal{I}_{K}$.
-We also store a set of node indices of top $N$ ranked functions $\mathcal{I}_{N}$ where $N>K$.
-This is our primary set of retrieved functions which we aim to further improve using neighborhood exploration.
-Neighborhood Exploration.
-Since we have a retrieval limit of $K$, it is not affordable to store the entire neighborhood of all the nodes in $\mathcal{I}_{K}$.
-To strictly limit the neighbors we explore and add to the final retrieval list of the same size $K$, we restrict the neighborhood exploration to the first $C$ nodes in $\mathcal{I}_{K}$ within a $d$-hop distance from the center nodes.
-While we primarily focus on exploring neighbors via the ‘contains’ edge type to capture the structural organization of the codebase, SpIDER can be extended to incorporate other edge types like ‘imports’, ‘invokes’, ‘inherits’, etc.
-for further performance enhancements.
-Two functions within the same class would be 2 hops away from each other.
-Similarly, two functions within two separate classes in two separate files will be 4 hops away.
-We do not explore neighbors beyond $d$ hop distance.
-Increasing the depth improves coverage at the cost of increased neighbors context per center.
-Hence, $d$ and $C$ are hyperparameters that a user can adjust based on the downstream budget, thereby offering more flexibility and tuning capabilities in terms of retrieval performance.
-For evaluation, we set $C=5$ and $N=500$ based on optimal performance observed for Python in the SWE-PolyBench dataset (see table˜4 and table˜12).
-These values are kept fixed across all languages and datasets to avoid hyperparameter overfitting.
-Additionally, we set $d=4$ to ensure that token consumption per LLM invocation remains below 10 00010\,000.
-While higher values of $K$ generally improve retrieval performance (see table˜13, fig.˜2), we evaluate for $K=20$ in our main results (Tables 3 and 2) to ensure fairness to the SweRank baseline, which selects $K=20$ as optimal for its downstream LLM-based reranker that improves localization at higher ranks.
-Neighbor filtering
-[…]
-6 Experiments
-Our goal is to answer the following three research hypotheses from our experiments:
-H1 Can the graph structure of a codebase be leveraged to improve retrieval recall and accuracy for instances involving multiple function edits that are spatially close within the code graph?
-H2 Among the top-$K$ nodes retrieved using dense embedding retrieval, do the highest-ranked nodes serve as accurate approximations of the actual ground truth locations in the code graph?
-If so, is it possible to reach the true nodes through local exploration of the code graph, potentially trading off some lower-ranked retrieved nodes in favor of better spatial coverage?
-H3 How do existing retrieval methods, along with our proposed approach, perform across programming languages beyond Python?
-To this end, we focus our study towards scenarios having more than 1 edit per instance as shown in table 1 for 4 programming languages, namely Python, Java, JavaScript and TypeScript, highlighting the robustness of the baselines and our approach.
+4.1 Building Code Graphs
+For a given codebase, we construct a code graph using only files written in the primary programming language, defined as the language comprising the majority of the repository.
+Files in other languages are ignored to simplify graph construction.
+For example, in a Python repository, we consider only Python source files.
+We follow the graph schema proposed by Chen et al.
+(2025), where each node represents a code entity—function, class, file, or directory—and edges encode structural relations of types imports, invokes, contains, and inherits.
+For Python, we use the built-in ast module [^1] https://docs.python.org/3/library/ast.html to parse source files and extract syntax trees.
+For Java, JavaScript, and TypeScript, we rely on Tree-sitter [^2] https://github.com/tree-sitter for syntax parsing.
+Graph construction proceeds by traversing the repository hierarchy.
+Directory nodes are added to enable navigation across the project structure.
+File nodes are included only if they contain at least one function or class definition; files containing solely metadata or configuration are ignored.
+Function and class nodes, along with the edges connecting them, are extracted from the corresponding syntax trees.
+While directory nodes do not store code content, all other node types include their associated source code.
+Constructing accurate graphs for Java, JavaScript and TypeScript presents additional challenges due to the flexible ways in which functions and classes can be declared/implemented.
+For example, class definitions and method implementations may be split across multiple locations or files.
+We explicitly handle such language-specific cases using Tree-sitter to ensure faithful graph construction.
+Refer App.
+A.8.
+Although the resulting graph contains multiple node types, we focus exclusively on retrieving functions, which constitute the most granular, and consequently the most challenging retrieval unit.
+This choice is motivated by statistics from the SWE-PolyBench dataset, where $67\%$ of instances involve function-level edits, with an average of $2.78$ edits per instance (Rashid et al., 2025).
+In contrast, only $1.42\%$ of instances involve class-only edits (excluding method changes), with an average of $0.49$ edits per instance.
+Since files are composed of functions and classes, accurate retrieval at finer granularities naturally helps file-level localization as well.
+Table 1 reports dataset statistics, including the proportion of retained instances per language.
+For function-level retrieval, we exclude instances involving only file- or class-level edits.
+Similarly, for class- and file-level retrieval, we ignore instances with edits exclusively at other granularities.
+4.2 Leveraging Code Graph via SpIDER
+fig.˜1 provides a high-level overview of SpIDER, with algorithmic details presented in algorithm˜1 (section˜A.7).
+Given an issue description $Q$ and a codebase represented as graph $\mathcal{G}$, SpIDER retrieves the top-$K$ functions most likely to require modification.
+While we focus on function-level retrieval, the framework naturally extends to class- and file-level retrieval by aggregating subsets of code chunks when full contexts exceed the LLM budget (see section˜A.9).
+SpIDER’s design is motivated by the observation that, in instances requiring multiple edits, relevant functions tend to be structurally proximate in the graph due to call dependencies and containment relationships.
+While dense semantic retrieval provides broad coverage, it often fails to capture “near-miss” cases, where the correct function lies close structurally but not semantically to top-ranked candidates.
+Semantic Retrieval  We first compute the baseline dense retrieval $\mathcal{S}_{K}(Q)$ as defined in section˜3.
+Additionally, we compute $\mathcal{S}_{N}(Q)$ containing the top-$N$ nodes ($N>K$) to constrain neighborhood exploration.
+Seed Selection  From $\mathcal{S}_{K}(Q)$, we select the top-$C$ ranked nodes as seed centers for graph exploration:
+$\mathcal{C}_{Q}:={v\in\mathcal{S}_{C}(Q)}$
+where $C\leq K$. A seed anchors a local neighborhood search.
+Neighborhood Exploration  For each seed, we perform breadth-first search up to depth $d$ along ‘contains’ edges, collecting structurally proximate function nodes.
+Formally, the $d$-hop neighborhood is:
+$\Gamma_{d}(\mathcal{C}_{Q}):=\{u\in\mathcal{V}\mid\mathrm{dist}_{\mathcal{G}}(u,\mathcal{C}_{Q})\leq d\}$
+where $\mathrm{dist}_{\mathcal{G}}(u,\mathcal{C}_{Q})$ denotes the shortest-path distance from $u$ to any node in $\mathcal{C}_{Q}$.
+Two functions within the same class are 2 hops apart; functions in different classes across separate files are 4 hops apart.
+While we use ‘contains’ edges to capture hierarchical structure, the framework can incorporate other edge types like ‘invokes’, ‘imports’, and ‘inherits’.
+Two-Stage Neighbor Filtering
+To control computational cost while maintaining quality, candidates are filtered in two stages.
+Stage 1 (Semantic Filtering): We restrict neighbors to those also appearing in the top-$N$ candidate pool:
+$\hat{\mathcal{C}}_{Q}:=\Gamma_{d}(\mathcal{C}_{Q})\cap\mathcal{S}_{N}(Q)$
+This ensures exploration remains within semantically relevant regions of the graph.
+Stage 2 (LLM-based Selection): An LLM selector $\mathcal{L}$ evaluates each candidate $u\in\hat{\mathcal{C}}_{Q}$, receiving the source code content and issue description as context.
+The LLM returns a binary relevance decision $\mathcal{L}(u)\in\{0,1\}$.
+Note that seed centers $\mathcal{C}_{Q}$ are always retained in $\mathcal{U}_{Q}$, where:
+$\mathcal{U}_{Q}:={u\in\hat{\mathcal{C}}_{Q}\mid\mathcal{L}(u)=1}\cup\mathcal{C}_{Q}$
+Output Construction  To maintain a fixed retrieval budget of $K$, newly selected neighbors replace the lowest-ranked non-center nodes. The final retrieval set is:
+$\mathcal{S}^{\texttt{LLM}}_{K}(Q):=\mathcal{U}_{Q}\cup\underset{v\in\mathcal{\mathcal{S}_{K}(Q)}}{\mathrm{arg\,top}_{K-|\mathcal{U}_{Q}|}}\;s_{Q}(v)$ (1)
+HyperparametersSpIDER introduces four hyperparameters: retrieval budget $K$, number of seed centers $C$, exploration depth $d$, and semantic filtering threshold $N$.
+We analyze their effects in section˜5.3.
+Definition 4.1 (Expected Recall Change).
+$\pi_{B}$ $:=\Pr(u\in\mathcal{V}^{*}\mid u\in\mathcal{S}_{K}\setminus\mathcal{S}_{K-|\mathcal{U}_{Q}|}),$
+$\pi_{\Gamma}$ $:=\Pr(u\in\mathcal{V}^{*}\mid u\in\hat{\mathcal{C}}_{Q}\setminus\mathcal{S}_{K}).$
+Proposition 4.2 (Sufficient Condition for Recall Improvement).
+Following the definitions from sections 3 and 4.2 and under the assumptions 1, 2 and 3 listed in section˜A.2, if it holds that $\alpha\pi_{\Gamma}>\pi_{B}+\beta(1-\pi_{\Gamma})$, then we can say,
+$\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}^{\texttt{LLM}}_{K})]>\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}_{K})].$
+Proposition 4.2 shows that when (i) relevant functions form a localized region in the code graph, (ii) dense retrieval places at least one seed in that region, and (iii) the LLM selector has higher true-positive than false-positive rate, then graph-aware retrieval strictly improves expected recall at fixed budget $K$.
+Proof can be found in section˜A.1.
+5 Experiments and Results
+We design our experiments to investigate whether leveraging structural relationships within code improves retrieval performance, particularly in challenging function-level settings.
+We focus on three high-level research questions:
+RQ1 (Exploiting Graph Structure) Can incorporating the structural relationships between code modules improve retrieval of relevant functions beyond what semantic similarity alone achieves?
+RQ2 (Cross-Language Robustness) Do retrieval strategies that leverage code structure generalize across programming languages and codebases of varying complexity?
+RQ3 (Efficiency and Practicality) Can graph-aware retrieval methods achieve improved coverage without incurring excessive computational cost or LLM invocations?
 5.1 Experimental Setup
-(a)  Python
-(b)  Java
-(c)  JavaScript
-(d)  TypeScript
-Figure 2: Dense retrieval methods performance comparison on SWE-PolyBench benchmark.
-Datasets
-We conduct experiments on SWEBench-Verified, SWE-PolyBench and Multi-SWEBench considering patched functions as the ground truth targets for retrieval.
-Following suresh2025cornstack, we exclude the instances having no function level modifications in their patch.
-We evaluate across the Python, Java, TypeScript, and JavaScript programming languages.
-Evaluation metrics
-Since, the goal is to improve the coverage when retrieving $K$ nodes, we mainly use Recall@$K$ and Acc@$K$ for evaluating and comparing various methods.
-Here Acc@$K$ for an instance is 1 if all the ground truth targets are present in $K$ retrieved predictions.
-We also present Mean Reciprocal Rank (MRR) which is the inverse of the rank of the first encounter of the ground truth node in the top-$K$ retrieved list.
-Hyperparameters
-The hyperparameters for our algorithm are $C$, $d$, NN and $K$.
-We present various studies ablating on each of these parameters in the subsequent sections.
-Following reddy2025sweranksoftwareissuelocalization; fehr-etal-2025-coret we use $K=20$ for all the datasets to be fair to localization methods that are sensitive to the choice of this hyperparameter such as SWERankEmbed + reranker.
-We perform a grid search to select SpIDER’s hyperparameters on SWE-PolyBench’s Python split, resulting in $C=5$, $d=4$ and $N=500$.
-We keep these values constant for all our experiments.
-In our ablation studies, we analyze how these design choices affect the LLM’s context length requirements and invocation budget, providing deeper insights for downstream users.
-For all experiments, we set the LLM temperature to $0.1$.
-Baselines
-We compare SpIDER against state-of-the-art retrieval approaches based on dense embedding models like SweRankEmbed-Small ZS, CodeSAGE-Small ZS, reddy2025sweranksoftwareissuelocalization; zhang2024codesage and BM25 Robertson1994OkapiATBM25 as it represent a popular sparse retrieval method.
-Because CodeSAGE-Small ZS is not trained for Github-style problem solving tasks, we further finetune it using parameter efficient finetuning using LoRA finetuning (hu2022lora) on $2000$ instances from 3 repositories of the SWEBench training dataset.
-We call this baseline CodeSAGE-Small-PEFT.
-fehr-etal-2025-coret did not release the embedding model as well as code publicly.
-Hence, we could not use it as a baseline.
-We do not include LocAgent (chen2025locagent) due to two main drawbacks: it is extremely slow and requires significantly more number of LLM invocations as opposed to SpIDER.
-To give an estimate of performance of LocAgent compared to other methods, we evaluate its performance on SWEBench-Verified and on Python repositories of SWE-PolyBench benchmark table˜9.
-It took us more than 50 hours to complete the evaluation of LocAgent on SWEBench-Verified while SpIDER finished the evaluation in less than 2 hours.
-Additionally, SpIDER requires $C$ invocations of an LLM which is not deterministic for LocAgent.
-We improve each baseline using spatial information from code graphs and demonstrate two retrieval approaches for each model, one using SpIDER and the other using dense retrieval with cosine similarity for dense embedding models and the standard sparse retrieval implementation for BM25.
+Figure 2: Function-level retrieval performance on SWE-PolyBench.
+Across all programming languages and retrieval budgets $K$, SWERankEmbed-Small+SpIDER consistently outperforms dense and sparse baselines in both Accuracy and Recall, highlighting robust gains from incorporating graph-aware neighborhood expansion under a fixed retrieval budget.
+Shaded envelopes indicate standard deviations estimated via bootstrapping with $1000$ draws.
+Datasets  To answer our research questions, we focus on challenging multi-function edits and evaluate retrieval performance across multiple programming languages.
+Specifically, we conduct experiments on SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench, using patched functions as the ground-truth targets for retrieval.
+Following Suresh et al.
+(2025), we exclude instances that do not contain function-level modifications.
+The evaluation spans Python, Java, TypeScript, and JavaScript repositories, allowing us to assess both the robustness of our approach and the generalization of baseline methods across different languages and codebase complexities.
+Evaluation Metrics  Given a query $Q$ and a retrieved set $\mathcal{S}(Q)$ of size $K$, we evaluate retrieval quality using Recall@$K$ and Acc@$K$, defined as
+$\text{Recall@}K=\frac{|\mathcal{V}^{*}\cap\mathcal{S}(Q)|}{|\mathcal{V}^{*}|},\qquad\text{Acc@}K=\mathbb{I}[\mathcal{V}^{*}\subseteq\mathcal{S}(Q)]$
+where $\mathcal{V}^{*}$ denotes the set of ground-truth patched functions.
+Recall@$K$ measures the fraction of ground-truth functions recovered within the top-$K$ retrieved results, while Acc@$K$ is a stricter metric that equals 1 only when all ground-truth functions are included in the top-$K$ set.
+These metrics are particularly well-suited to the multi-function setting, as they directly capture coverage under a fixed retrieval budget.
+For completeness, we additionally report Mean Reciprocal Rank (MRR) in App.
+section˜A.10, defined as the reciprocal of the rank at which any ground-truth function first appears in the retrieved list.
+Hyperparameters  We analyze the effects of hyperparameters $C$, $d$, NN, and $K$ in section˜5.3.
+Following prior work Reddy et al.
+(2025); Fehr et al.
+(2025), we fix the retrieval budget to $K=20$ across all datasets to ensure fair comparison with baselines that are sensitive to this choice, such as embedding-based retrieval followed by LLM reranking.
+We select the remaining hyperparameters of SpIDER via grid search on the Python split of SWE-PolyBench, resulting in $C=5$, $d=4$, and $N=500$.
+These values are held constant for all experiments to avoid overfitting.
+In our ablation studies, we further examine how varying these parameters impacts retrieval performance as well as practical considerations such as LLM context length requirements and invocation budget.
+For all experiments, we use Claude Sonnet 4 with a temperature of $0.1$ to ensure consistent and reproducible evaluation.
+Baselines  We compare SpIDER against representative state-of-the-art retrieval approaches spanning both dense and sparse methods.
+For dense retrieval, we consider SweRankEmbed-Small ZS and CodeSAGE-Small ZS, which are widely used embedding-based models for software issue localization and code retrieval Reddy et al.
+(2025); Zhang et al.
+(2024).
+As a sparse retrieval baseline, we include BM25 Robertson et al.
+(1994).
+Since CodeSAGE-Small ZS is not trained for GitHub-style problem-solving tasks, we additionally evaluate a parameter-efficiently fine-tuned variant using LoRA (Hu et al., 2022) on $2000$ instances from three repositories in the SWEBench training set, denoted as CodeSAGE-Small-PEFT.
+Some recent methods could not be fully evaluated.
+Fehr et al.
+(2025) did not release their embedding model or code, preventing direct comparison.
+LocAgent (Chen et al., 2025) is based on a time consuming and iterative LLM-driven exploration and is therefore evaluated only on SWEBench-Verified and the Python repositories of SWE-PolyBench in tables 8 and 9.
 5.2 Results
-Table 2: Comparison of dense retrieval and dense retrieval + reranker/selector for $K=20$ and $N=500$, $C=5$, $d=4$. Winners per metric are highlighted in light blue.
-Language Model Retrieval Dense Retrieval Only Dense + Reranker/Selector
-Recall@20 Accuracy@20 MRR@20 Recall@3 Accuracy@3 MRR@3
-Python SweRankEmbed-Small ZS SpIDER 0.49 0.40 0.30 0.44 0.36 0.54
-DER 0.42 0.35 0.26 0.38 0.31 0.48
-CodeSAGE-Small ZS SpIDER 0.27 0.21 0.13 0.23 0.18 0.31
-DER 0.21 0.17 0.11 0.19 0.16 0.25
-CodeSAGE-Small-PEFT SpIDER 0.37 0.27 0.24 0.32 0.25 0.42
-DER 0.29 0.20 0.21 0.26 0.20 0.38
-BM25 SpISR 0.33 0.26 0.19 0.30 0.23 0.38
-SR 0.31 0.24 0.18 0.27 0.22 0.35
-Java SweRankEmbed-Small ZS SpIDER 0.36 0.27 0.18 0.25 0.17 0.36
-DER 0.31 0.24 0.18 0.24 0.16 0.33
-CodeSAGE-Small ZS SpIDER 0.23 0.17 0.10 0.16 0.09 0.26
-DER 0.15 0.10 0.07 0.11 0.07 0.19
-CodeSAGE-Small-PEFT SpIDER 0.28 0.18 0.15 0.20 0.14 0.28
-DER 0.21 0.14 0.13 0.17 0.11 0.25
-BM25 SpISR 0.16 0.10 0.10 0.12 0.08 0.20
-SR 0.15 0.10 0.10 0.12 0.09 0.19
-JavaScript SweRankEmbed-Small ZS SpIDER 0.60 0.52 0.18 0.43 0.37 0.43
-DER 0.52 0.45 0.16 0.41 0.34 0.42
-CodeSAGE-Small ZS SpIDER 0.38 0.32 0.09 0.30 0.25 0.31
-DER 0.26 0.21 0.07 0.22 0.17 0.24
-CodeSAGE-Small-PEFT SpIDER 0.55 0.48 0.15 0.41 0.35 0.41
-DER 0.45 0.38 0.14 0.36 0.30 0.39
-BM25 SpISR 0.46 0.39 0.13 0.36 0.30 0.38
-SR 0.36 0.29 0.11 0.31 0.25 0.34
-TypeScript SweRankEmbed-Small ZS SpIDER 0.35 0.28 0.06 0.27 0.21 0.29
-DER 0.29 0.21 0.18 0.25 0.19 0.30
-CodeSAGE-Small ZS SpIDER 0.16 0.10 0.03 0.14 0.09 0.19
-DER 0.10 0.06 0.06 0.09 0.06 0.14
-CodeSAGE-Small-PEFT SpIDER 0.25 0.18 0.04 0.22 0.16 0.27
-DER 0.19 0.13 0.11 0.18 0.12 0.24
-BM25 SpISR 0.11 0.09 0.02 0.10 0.07 0.11
-SR 0.07 0.05 0.04 0.07 0.05 0.09
-Leveraging Code Graphs Improves Retrieval
-We assess the effect of graph structure on code retrieval in SWE-PolyBench, comparing standard dense embeddings to our graph-aware SpIDER across Python, TypeScript, JavaScript, and Java (Fig.
-2).
-The top-$K$ modules are retrieved by ranking code embeddings against issue embeddings via cosine similarity.
-SpIDER (SweRankEmbed-Small + SpIDER), which incorporates spatially informed graph structure, consistently outperforms all baselines across languages and $K$ values, highlighting the benefit of leveraging codebase graph information.
-Notably, the zero-shot SweRankEmbed-Small model– trained solely on $\approx$3,300 Python repositories– remains competitive across non-Python languages, surpassing both BM25 and codesage-small-v2 in zero-shot and PEFT settings, demonstrating strong cross-lingual transfer.
-Overall, while semantic embeddings provide a solid foundation, explicitly modeling spatial relationships via SpIDER yields substantial additional gains in retrieval accuracy.
-Table 3: Dense retrieval performance across benchmarks for $K=20$, $N=500$, $C=5$, $d=4$. Winners per metric highlighted in light blue.
-Language Model Retrieval Recall@20 Accuracy@20 MRR@20
+Table 2: Main result.
+Across all embedding models and programming languages, SpIDER consistently improves Recall and Accuracy over dense retrieval alone, with gains preserved and often amplified after LLM reranking.
+Here, $K=20$ with $N=500$, $C=5$, and $d=4$; best results per embedding model are highlighted in light blue.
+SpISR denotes the spatially informed sparse retriever analogous to SpIDER.
+KDE plots are provided in Figs.
+4 and 5 while confidence intervals are provided in table˜10.
+Language Model Retrieval SWE-PolyBench SWEBench-Verified
+Dense Retrieval Only Dense + LLM reranker Dense Retrieval Only
+Recall@20 Accuracy@20 Recall@3 Accuracy@3 Recall@20 Accuracy@20
+Python SweRankEmbed-Small ZS SpIDER 0.49 0.40 0.44 0.36 0.61 0.56
+DER 0.42 0.35 0.38 0.31 0.54 0.49
+CodeSAGE-Small ZS SpIDER 0.27 0.21 0.23 0.18 0.30 0.27
+DER 0.21 0.17 0.19 0.16 0.21 0.19
+CodeSAGE-Small-PEFT SpIDER 0.37 0.27 0.32 0.25 0.55 0.49
+DER 0.29 0.20 0.26 0.20 0.43 0.38
+BM25 SpISR 0.33 0.26 0.30 0.23 0.45 0.41
+SR 0.31 0.24 0.27 0.22 0.38 0.34
 Multi-SWEBench
-Java SweRankEmbed-Small ZS SpIDER 0.37 0.28 0.08
-DER 0.32 0.24 0.08
-CodeSAGE-Small ZS SpIDER 0.21 0.19 0.05
-DER 0.15 0.13 0.03
-CodeSAGE-Small-PEFT SpIDER 0.31 0.25 0.10
-DER 0.23 0.17 0.09
-BM25 SpISR 0.23 0.17 0.07
-SR 0.22 0.17 0.05
-JavaScript SweRankEmbed-Small ZS SpIDER 0.39 0.30 0.18
-DER 0.31 0.23 0.15
-CodeSAGE-Small ZS SpIDER 0.17 0.12 0.08
-DER 0.09 0.06 0.05
-CodeSAGE-Small-PEFT SpIDER 0.32 0.24 0.13
-DER 0.25 0.18 0.11
-BM25 SpISR 0.21 0.14 0.08
-SR 0.14 0.10 0.06
-TypeScript SweRankEmbed-Small ZS SpIDER 0.52 0.42 0.08
-DER 0.40 0.32 0.23
-CodeSAGE-Small ZS SpIDER 0.33 0.26 0.04
-DER 0.26 0.21 0.13
-CodeSAGE-Small-PEFT SpIDER 0.30 0.26 0.05
-DER 0.19 0.15 0.12
-BM25 SpISR 0.32 0.27 0.04
-SR 0.24 0.21 0.11
-SWEBench-Verified
-Python SweRankEmbed-Small ZS SpIDER 0.61 0.56 0.32
-DER 0.54 0.49 0.29
-CodeSAGE-Small ZS SpIDER 0.30 0.27 0.13
-DER 0.21 0.19 0.10
-CodeSAGE-Small-PEFT SpIDER 0.55 0.49 0.27
-DER 0.43 0.38 0.22
-BM25 SpISR 0.45 0.41 0.25
-SR 0.38 0.34 0.22
-Improvement is Consistent Across Datasets, Metrics, % Languages
-Next, we augment all baseline retrieval methods—including dense models (SweRankEmbed-Small, codesage-small-v2) and BM25 sparse retrievers– with code graph structure information and measure the resulting improvements.
-We perform this analysis across all datasets (SWEBench-Verified, SWE-PolyBench, Multi-SWEBench) as reported in Tab.
-2 and 3.
-Our results reveal that our SpIDER strategy consistently enhances retrieval performance, particularly in Recall@K and Acc@K, across all programming languages and datasets.
-The best-performing embedding model across datasets is SweRankEmbed-Small, and applying SpIDER further increases Recall@20 and Acc@20 by at least $12.9\%$ and $14\%$, respectively.
+Java SweRankEmbed-Small ZS SpIDER 0.36 0.27 0.25 0.17 0.37 0.28
+DER 0.31 0.24 0.24 0.16 0.32 0.24
+CodeSAGE-Small ZS SpIDER 0.23 0.17 0.16 0.09 0.21 0.19
+DER 0.15 0.10 0.11 0.07 0.15 0.13
+CodeSAGE-Small-PEFT SpIDER 0.28 0.18 0.20 0.14 0.31 0.25
+DER 0.21 0.14 0.17 0.11 0.23 0.17
+BM25 SpISR 0.16 0.10 0.12 0.08 0.23 0.17
+SR 0.15 0.10 0.12 0.09 0.22 0.17
+JavaScript SweRankEmbed-Small ZS SpIDER 0.60 0.52 0.43 0.37 0.39 0.30
+DER 0.52 0.45 0.41 0.34 0.31 0.23
+CodeSAGE-Small ZS SpIDER 0.38 0.32 0.30 0.25 0.17 0.12
+DER 0.26 0.21 0.22 0.17 0.09 0.06
+CodeSAGE-Small-PEFT SpIDER 0.55 0.48 0.41 0.35 0.32 0.24
+DER 0.45 0.38 0.36 0.30 0.25 0.18
+BM25 SpISR 0.46 0.39 0.36 0.30 0.21 0.14
+SR 0.36 0.29 0.31 0.25 0.14 0.10
+TypeScript SweRankEmbed-Small ZS SpIDER 0.35 0.28 0.27 0.21 0.52 0.42
+DER 0.29 0.21 0.25 0.19 0.40 0.32
+CodeSAGE-Small ZS SpIDER 0.16 0.10 0.14 0.09 0.33 0.26
+DER 0.10 0.06 0.09 0.06 0.26 0.21
+CodeSAGE-Small-PEFT SpIDER 0.25 0.18 0.22 0.16 0.30 0.26
+DER 0.19 0.13 0.18 0.12 0.19 0.15
+BM25 SpISR 0.11 0.09 0.10 0.07 0.32 0.27
+SR 0.07 0.05 0.07 0.05 0.24 0.21
+Table 3: Ablation on number of seed centers $C$ using SweRankEmbed-Small+SpIDER. Increasing $C$ improves Recall@20 and Accuracy@20 across languages up to saturation.
+$\boldsymbol{C}$ Python JavaScript Java TypeScript
+Recall@20 Acc@20 Recall@20 Acc@20 Recall@20 Acc@20 Recall@20 Acc@20
+1 0.47 0.38 0.56 0.48 0.32 0.24 0.32 0.24
+3 0.49 0.39 0.59 0.51 0.35 0.26 0.31 0.24
+5 0.49 0.40 0.60 0.52 0.36 0.27 0.35 0.28
+7 0.49 0.39 0.60 0.52 0.36 0.27 0.37 0.30
+9 0.48 0.39 0.60 0.53 0.37 0.27 0.39 0.31
+Table 4: Ablation on BFS exploration depth ($d$) using SpIDER + SweRankEmbed-Small for $K=20$, $N=100$, $C=3$.
+Increasing $d$ improves Recall@20 and Accuracy@20 across languages, while LLM input/output tokens scale with $d$.
+$\boldsymbol{d}$ Python JavaScript Java TypeScript
+Recall@20 Acc@20 Tokens Recall@20 Acc@20 Tokens Recall@20 Acc@20 Tokens Recall@20 Acc@20 Tokens
+Input Output Input Output Input Output Input Output
+2 0.45 0.36 2.6K 32 0.54 0.46 2.0K 50 0.33 0.25 2.2K 22 0.31 0.23 1.5K 21
+4 0.46 0.37 3.6K 35 0.57 0.49 3.1K 25 0.33 0.25 2.1K 54 0.30 0.24 2.0K 24
+6 0.50 0.41 11.9K 48 0.62 0.53 10.1K 42 0.33 0.25 3.6K 67 0.36 0.28 4.3K 33
+8 0.55 0.44 29.7K 69 0.67 0.58 15.0K 48 0.34 0.27 5.6K 82 0.36 0.28 5.6K 37
+Leveraging code graphs improves retrievaltable˜2 demonstrates the benefit of incorporating code graph structure into retrieval.
+Comparing standard dense (DER) and sparse (SR) retrieval with our graph-aware variant (SpIDER) across Python, TypeScript, JavaScript, and Java repositories from SWE-PolyBench, Multi-SWEBench, and SWEBench-Verified at $K=20$, we observe consistent improvements across all datasets and languages.
 Fig.
-6 shows that SpIDER consistently outperforms baseline DER methods across varying numbers of function-level edits when using SweRankEmbed-Small.
-With the LLM temperature set to 0.1, the standard deviation across multiple iterations was $\leq 0.01$ for all metrics, indicating stable performance.
-MRR@$K$ Gains Depend on Graph Neighborhoods
-The MRR@$K$ metric benefits from SpIDER primarily when the first relevant ground-truth item within the top-$K$ results is a neighbor of a center node.
-In cases where the first relevant item is a non-center node, particularly in TypeScript as shown in Tab.
-3SpIDER may rank neighbors of center nodes ahead, slightly reducing MRR@$K$.
-Since SpIDER is not explicitly designed to optimize the precise ranking order within the top-$K$ results, improvements in MRR@$K$ over baseline DER methods are modest but still observable.
-Retrieval Strategy is Agnostic to Downstream Rerankers
-Our retrieval strategy is agnostic to downstream rerankers or localization steps.
-To validate this, we evaluate an LLM selector (similar to the reranker in reddy2025sweranksoftwareissuelocalization) that retrieves the top 3 modules from a set of $K=20$ candidates.
-As reported in Tab.
-2, improving top-$K$ coverage with SpIDER also translates to notable gains in Recall@20 and Acc@20 for top-3 localized modules.
+2 further illustrates this trend on SWE-PolyBench as the retrieval budget $K$ varies.
+SweRankEmbed-Small + SpIDER consistently outperforms all baselines across languages and $K$ values, highlighting the effectiveness of combining semantic similarity with explicit modeling of spatial relationships and LLM-based neighborhood filtering.
+Notably, the zero-shot SweRankEmbed-Small model—trained exclusively on roughly 3,300 Python repositories—remains competitive on non-Python codebases, outperforming BM25 and codesage-small-v2 in both zero-shot and PEFT settings, demonstrating strong cross-language transfer.
+Across datasets, SpIDER improves Recall@20 and Acc@20 of SweRankEmbed-Small by at least $13\%$ and $14\%$, respectively, indicating that graph-aware retrieval yields substantial performance gains even when paired with already strong semantic embeddings.
+Improvement is consistent across datasets, metrics, and languages  As shown in fig.˜3, SpIDER yields consistent improvements across both context configurations, achieving gains of +0.46+0.46 percentage points (pp) at $K=5$ and +2.74+2.74 pp at $K=10$, corresponding to 2 and 12 additional resolved instances, respectively.
+These results demonstrate that enhanced retrieval coverage directly translates into measurable gains in downstream code generation performance, completing the chain from graph-aware retrieval to end-task success.
+Retrieval strategy is agnostic to downstream rerankers  Our retrieval strategy is designed to operate independently of downstream reranking or localization components.
+To validate this property, we evaluate all retrieval methods in conjunction with an LLM-based reranker following Reddy et al.
+(2025), which selects the top-3 functions from $K=20$ retrieved candidates based on its reasoning capabilities.
+As shown in table˜2, the improved top-20 coverage provided by SpIDER consistently translates into higher Recall@3 and Acc@3 after reranking.
+This implies that gains achieved at the retrieval stage are preserved and often amplified by downstream reranking, and that SpIDER is complementary to, rather than coupled with, specific localization strategies.
+Improved retrieval translates to better code generation  To evaluate the practical impact of improved retrieval on downstream task performance, we pass the retrieved function paths to swe-mini-agent Yang et al.
+(2024) executed with Claude Sonnet 4.5, measuring pass rate on the SWE-Bench-Verified benchmark with retrieval budgets $K=5$ and $K=10$.
+As shown in fig.˜3, SpIDER yields consistent improvements across both context configurations, achieving gains of +0.46+0.46 percentage points (pp) at $K=5$ and +2.74+2.74 pp at $K=10$, corresponding to 2 and 12 additional resolved instances, respectively.
+These results demonstrate that enhanced retrieval coverage directly translates into measurable gains in downstream code generation performance, completing the chain from graph-aware retrieval to end-task success.
+Figure 3: Pass rate improvement of SWERankEmbed-Small-SpIDER over SWERank on SWE-bench Verified with $C=5,d=4,N=500$.
 5.3 Ablation Experiments
-Table 4: Impact of the number of center nodes ($C$) on retrieval performance using SpIDER + SweRankEmbed-Small.
-Increasing $C$ generally improves Recall@20 and Acc@20 across languages, while MRR@20 shows modest variation.
-$\boldsymbol{C}$ Python JavaScript Java TypeScript
-Recall@20 Acc@20 MRR@20 Recall@20 Acc@20 MRR@20 Recall@20 Acc@20 MRR@20 Recall@20 Acc@20 MRR@20
-1 0.47 0.38 0.29 0.56 0.48 0.18 0.32 0.24 0.18 0.32 0.24 0.06
-3 0.49 0.39 0.30 0.59 0.51 0.18 0.35 0.26 0.19 0.31 0.24 0.06
-5 0.49 0.40 0.30 0.60 0.52 0.18 0.36 0.27 0.18 0.35 0.28 0.06
-7 0.49 0.39 0.30 0.60 0.52 0.18 0.36 0.27 0.19 0.37 0.30 0.06
-9 0.48 0.39 0.30 0.60 0.53 0.19 0.37 0.27 0.19 0.39 0.31 0.06
-Table 5: Impact of the number of neighbors per center ($d$) on retrieval performance and LLM token usage using SpIDER + SweRankEmbed-Small for $K=20$, $N=100$, $C=3$.
-Increasing $d$ improves Recall@20 and Accuracy@20 across languages, while LLM input/output tokens scale with $d$.
-Best values per metric are highlighted.
-$\boldsymbol{d}$ Python JavaScript Java TypeScript
-Recall@20 Acc@20 Input Output Recall@20 Acc@20 Input Output Recall@20 Acc@20 Input Output Recall@20 Acc@20 Input Output
-2 0.45 0.36 2,642 32 0.54 0.46 2,032 50 0.33 0.25 2,158 22 0.31 0.23 1,474 21
-4 0.46 0.37 3,591 35 0.57 0.49 3,115 25 0.33 0.25 2,083 54 0.30 0.24 1,994 24
-6 0.50 0.41 11,904 48 0.62 0.53 10,065 42 0.33 0.25 3,615 67 0.36 0.28 4,334 33
-8 0.55 0.44 29,658 69 0.67 0.58 15,028 48 0.34 0.27 5,594 82 0.36 0.28 5,643 37
-On the Number of Top $C$ Centers
-In this section, we provide ablation studies on varying the number of top $C$ centers on SWE-PolyBench for retrieval performance.
-Increasing $C$ results in more neighborhood explorations and addition of relevant neighbors in the list of top-$K$ retrieved modules at the cost of dropping bottom semantically ranked modules.
-Furthermore, each neighborhood exploration around a center requires 1 LLM call for filtering neighbors.
-Hence, increasing $C$ results in increased number of LLM calls.
-In table˜4, we ablate on $C$ and observe the corresponding performance changes.
-We see that the performance saturates at $C=5$ for $N=500,d=4$ at $K=20$.
-Hence, we stick with $C=5$ for all our experiments.
-A downstream user can aim to choose a $C$ that fits in their LLM invocation limits per instance.
-On $K$
-The choice of $K$ for retrieval depends on the budget of the downstream application.
-Retrieval performance is expected to improve as $K$ increases.
-We further validate this claim in figures 2 and 5 where we see a monotonic improvement in performance as $K$ incremented.
-The downstream user can choose a $K$ that is allowed by their budget while aiming for best retrieval performance.
-On depth $d$
-In this section, we vary the maximum depth $d$ for breadth-first search (BFS) exploration while selecting neighbors and observe the corresponding impact on the performance.
-Greater depth would result in more number of neighbors per center that need to be filtered using the two-stage neighborhood filtering.
-In table˜5, we ablate on $d$ keeping $N=100,C=3$ and $K=20$.
-We see that there is notable improvement in the Recall@20 and Acc@20 performance with increasing depth $d$, however, increasing $d$ also increases the token consumption to LLM while filtering the neighbors.
-The downstream user can choose $d$ based on their token budget per LLM invocation.
-On the Primary Neighborhood Filtering Threshold NN
-In table˜12, ablate on $N$ to see it’s impact on the retrieval performance.
-Higher $N$ results in more neighbors passing the primary filteration and hence, more neighbors for the LLMs to sift from.
-We see that the performance saturates around $N=300$ for $C=3,K=20,d=4$.
-Hence, the choice of $N$ offers the same flexibility as $d$ to the downstream user.
-The user can choose $N$ based on the token limit per LLM call.
+To better understand how SpIDER’s hyperparameters influence retrieval performance and to provide actionable guidance for downstream users, we conduct ablation studies on four key parameters: number of seed centers $C$, retrieval budget $K$, BFS exploration depth $d$ and primary neighborhood filtering threshold $N$.
+For each hyperparameter, we follow a consistent structure: we first describe the motivation for the parameter, then its effect on retrieval performance, discuss trade-offs in terms of computational cost/LLM usage, and finally provide guidance for tuning in practice.
+These experiments aim to help users configure SpIDER effectively for different codebases and resource budgets.
+Number of Top $C$ Centers  The number of seed centers $C$ determines how many top-ranked nodes are used as starting points for neighborhood exploration.
+Increasing $C$ expands the explored neighborhood, adding more relevant neighbors to the top-$K$ retrieved nodes but potentially displacing lower-ranked semantic candidates.
+Each neighborhood exploration requires one LLM call, so higher $C$ increases total LLM invocations.
+table˜3 shows that performance saturates at $C=5$ for $N=500$, $d=4$, and $K=20$, indicating that exploring neighbors of the top-ranked nodes efficiently recovers ground-truth functions while keeping LLM usage reasonable.
+Users can adjust $C$ according to their LLM budget.
+Retrieval Budget $K$  The retrieval budget $K$ controls the number of candidate nodes returned.
+As expected, performance improves monotonically with increasing $K$ (figs.˜2 and 6), since a larger pool increases the chance of including relevant nodes.
+Users can select $K$ based on the constraints of downstream tasks, balancing retrieval quality against processing cost.
+Exploration Depth $d$  The maximum BFS depth $d$ determines how many neighbors are explored around each seed center.
+Larger $d$ increases the number of candidate neighbors, improving Recall@20 and Acc@20 (table˜4) but also increasing token consumption during the LLM-based neighborhood filtering.
+Users should select $d$ considering the trade-off between retrieval gains and LLM resource usage.
+Primary Neighborhood Filtering Threshold NN  The threshold $N$ controls how many neighbors pass the initial semantic filter before LLM-based refinement.
+Higher $N$ allows more neighbors to be considered, increasing the chance of recovering relevant nodes but also consuming more LLM tokens.
+Ablations in table˜11 show performance saturates around $N=300$ for $C=3$, $d=4$, and $K=20$.
+Like $d$, NN can be tuned based on downstream LLM token budgets.
+Takeaways  Overall, these ablations show that SpIDER’s hyperparameters offer a clear trade-off between retrieval performance and LLM cost.
+Key settings i.e., $C=5$, $d=4$, $N=500$, and $K=20$ provide strong performance across SWE-PolyBench, while users can tune $C$, $d$, NN, and $K$ according to resource constraints.
+Our results highlight that carefully leveraging top-ranked nodes and their neighborhoods is an effective strategy for maximizing retrieval coverage while controlling computational cost.
 6 Conclusion
@@
 Appendix A Appendix
-A.1 Graph construction details
+A.1 Theoretical Analysis
+The proof for proposition˜4.2 is simple.
+We start by providing a theoretical justification for why incorporating graph locality and LLM-based neighbor selection improves retrieval performance over standard dense embedding retrieval.
+Our analysis focuses on recall at a fixed budget $K$ and relies on mild structural and statistical assumptions.
+A.2 Modeling Assumptions
+We follow the problem definition from section˜3 and use $\mathcal{S}_{K}$ to denote $\mathcal{S}_{K}(Q)$ for notational convenience. Then we make the following assumptions.
+Assumption 1 (Score Distributions).
+There exist distributions $F_{1}$ and $F_{0}$ on $\mathbb{R}$ such that
+$s_{Q}(v)\sim\begin{cases}F_{1},&v\in\mathcal{V}^{*},\\ F_{0},&v\notin\mathcal{V}^{*},\end{cases}$
+where $F_{i}(t)=\Pr[s_{Q}(v)\leq t\mid s_{Q}(v)\sim F_{i}]$ and $F_{1}$ first-order stochastically dominates $F_{0}$, i.e.,
+$F_{1}(t)\leq F_{0}(t)\quad\forall t\in\mathbb{R}.$
+(See Shaked and Shanthikumar [2007] for background on stochastic dominance.)
+Assumption 2 (Graph Locality of Relevance).
+There exists an integer $d\geq 1$ such that for any $v\in\mathcal{V}^{*}$,
+$\Pr\big(u\in\mathcal{V}^{*}\mid u\in\Gamma_{d}(\{v\})\big)\geq\Pr(u\in\mathcal{V}^{*}).$
+That is, relevance is non-negatively correlated with graph proximity.
+Assumption 3 (LLM Selector Accuracy).
+For any $u\in\mathcal{C}_{Q}$,
+$\Pr(\mathcal{L}(u)=1\mid u\in\mathcal{V}^{*})$ $=\alpha$
+$\Pr(\mathcal{L}(u)=1\mid u\notin\mathcal{V}^{*})$ $=\beta,$
+where $0\leq\beta<\alpha\leq 1$.
+A.3 Expected Recall Change
+Define
+$B_{Q}$ $=\mathcal{S}_{K}\setminus\mathcal{S}_{K-|\mathcal{U}_{Q}|}$
+$\pi_{B}$ $:=\Pr(u\in\mathcal{V}^{*}\mid u\in B_{Q}),$
+$\pi_{\Gamma}$ $:=\Pr(u\in\mathcal{V}^{*}\mid u\in\hat{\mathcal{C}}_{Q}\setminus\mathcal{S}_{K}).$
+Here, $B_{Q}$ refers to bottom $|\mathcal{U}_{Q}|$ ranked nodes in $\mathcal{S}_{K}(Q)$.
+Proposition A.1 (Expected Recall Difference).
+Under Assumptions 1–3, the expected change in recall satisfies
+$\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}^{\texttt{LLM}}_{K})]-\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}_{K})]=\frac{\alpha\,\mathbb{E}[|\mathcal{U}_{Q}\cap\mathcal{V}^{*}|]-\mathbb{E}[|B_{Q}\cap\mathcal{V}^{*}|]-\beta\,\mathbb{E}[|\mathcal{U}_{Q}\setminus\mathcal{V}^{*}|]}{|\mathcal{V}^{*}|}.$
+Proof.
+By construction,
+$|\mathcal{S}^{\texttt{LLM}}_{K}\cap\mathcal{V}^{*}|=|\mathcal{S}_{K}\cap\mathcal{V}^{*}|-|B_{Q}\cap\mathcal{V}^{*}|+|\mathcal{U}_{Q}\cap\mathcal{V}^{*}|.$
+Taking expectations and dividing by $|\mathcal{V}^{*}|$ yields the result. ∎
+A.4 Sufficient Condition for Improvement
+Proposition A.2 (Restating Sufficient Condition for Recall Improvement).
+Under the assumptions 1, 2 and 3 listed in section˜A.2, if
+$\alpha\pi_{\Gamma}>\pi_{B}+\beta(1-\pi_{\Gamma})$, then it holds that
+$\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}^{\texttt{LLM}}_{K})]>\mathbb{E}[\mathrm{Rec}@K(\mathcal{S}_{K})].$
+Proof.
+By linearity of expectation,
+$\mathbb{E}[|\mathcal{U}_{Q}\cap\mathcal{V}^{*}|]=\alpha K\pi_{\Gamma},\quad\mathbb{E}[|\mathcal{U}_{Q}\setminus\mathcal{V}^{*}|]=K(1-\pi_{\Gamma}).$
+Similarly,
+$\mathbb{E}[|B_{Q}\cap\mathcal{V}^{*}|]=K\pi_{B}.$
+Substituting into Proposition A.1 and simplifying yields the stated claim. This concludes the proof for proposition˜4.2
+A.5 Graph-Structured Relevance
+We now show that $\pi_{\Gamma}$ is strictly larger than $\pi_{B}$ under a standard cluster assumption.
+Definition (Relevant Subgraph).
+Let $H_{q}=(V_{q},E_{q})$ be an induced subgraph such that $\mathcal{V}^{*}\subseteq V_{q}$ and
+$\operatorname{diam}(H_{q})\leq d.$
+Proposition A.3 (Graph Expansion Enriches Relevance).
+Suppose:
+  1. 1.
+    There exists at least one seed $v\in C_{q}\cap V_{q}$.
+  2. 2.
+    Assumptions 1 and 2 hold.
+Then
+$\pi_{\Gamma}\geq\Pr(u\in\mathcal{V}^{*}),$
+with strict inequality unless $\mathcal{V}^{*}=V_{q}$.
+Proof.
+Since $\operatorname{diam}(H_{q})\leq d$, all nodes in $V_{q}$ lie in $\Gamma_{d}(\{v\})$.
+By Assumption 2, conditioning on membership in $\Gamma_{d}(\{v\})$ increases the probability of relevance.
+Restricting further to $\operatorname{TopN}$ preserves this enrichment by Assumption 1 (stochastic dominance).
+∎
+A.6 Conclusion
+The above results show that when (i) relevant functions form a localized region in the code graph, (ii) dense retrieval places at least one seed in that region, and (iii) the LLM selector has higher true-positive than false-positive rate, then graph-aware retrieval strictly improves expected recall at fixed budget $K$.
+A.7 Algorithm
+Algorithm 1 SpIDER
+0: Code graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, issue description $Q$, bi-modal encoder $\mathcal{F}(\cdot)$
+0: Parameters: retrieval budget $K$, filtering threshold NN, number of centers $C$, search depth $d$
+1:// Semantic Retrieval
+2: Compute node embeddings: $\{\mathcal{F}(v)\}_{v\in\mathcal{V}}$
+3: Compute query embedding: $\mathcal{F}(Q)$
+4: Compute relevance scores: $s_{Q}(v)=\cos(\mathcal{F}(v),\mathcal{F}(Q))$ for all $v\in\mathcal{V}$
+[…]
+[…]
+7:// Seed Selection
+[…]
+9:// Neighborhood Exploration
+[…]
+11:// Two-Stage Neighbor Filtering
+[…]
+[…]
+14:// Output Construction
+[…]
+16:return $\mathcal{S}(Q)$
+Note: The LLM selector $\mathcal{L}$ receives the source code content of all candidates $u\in\hat{\mathcal{C}}_{Q}$ for a given center, along with the issue description $Q$.
+It returns $\mathcal{L}(u)=1$ for candidates deemed relevant to resolving the issue, and 0 otherwise.
+A.8 Graph construction details
 JavaScript did not have a unique class data type until 2015 as it was only introduced ECMAScript 2015 also known as ES6.
@@
 a nested function with the definition of container function at some other location in the same/different file).
-A.2 Extension to Class-level edits
+A.9 Extension to Class-level edits
 Table 5: Class-level performance on SWE-PolyBench for $K=20$, $N=500$, $C=5$, $d=2$. Winners per metric are in bold.
@@
 SR 0.33 0.26 0.25
-TypeScript SweRankEmbed-Small ZS SpIDER
-DER
-CodeSAGE-Small ZS SpIDER
-DER
-CodeSAGE-Small-PEFT SpIDER
-DER
-BM25 SpISR
-SR
-In table˜6, we demonstrate the efficacy of the proposed approach at a class level.
+In table˜5, we demonstrate the efficacy of the proposed approach at a class level.
 Here, we include only Python and Java here, excluding JavaScript, as most JavaScript repositories in the SWE-PolyBench dataset lack significant use of class syntax (introduced in ECMAScript 2015), resulting in very few instances with class ground truth nodes.
 A ground truth edit is classified as a class-level edit (and the corresponding class as a ground truth class node), if the patch introduces a modification within any part of the class declaration or definition (e.g., constructors, destructors, or methods) to address the issue description.
-A.3 Additional Experiments
+A.10 Additional Experiments
 Figure 4: KDE (Kernel Density Estimate) plots with bootstrapped results of SWE-PolyBench benchmark for Recall@20 performance across various retrieval methods.
@@
 1000 0.48 0.39 0.31 0.59 0.51 0.18 0.36 0.27 0.19 0.32 0.24 0.06
-Table 9: LocAgent performance on SWEBench-Verified and the Python-only subset of SWE-PolyBench.
+Table 8: LocAgent performance on SWEBench-Verified and a subset of Python instances in SWE-PolyBench.
 Running LocAgent took over 48 hours on SWEBench-Verified and approximately 24 hours on SWE-PolyBench, highlighting a key limitation of this approach.
@@
 (Python subset) MRR@K 0.21 0.23 0.25 0.25 0.25 0.26 0.26 0.26 0.26 0.26 0.26 0.26
-Table 10: Comparison of LocAgent with dense retrieval methods + reranker for Python repositories in SWE-PolyBench and SWEBench-Verified
+Table 9: Comparison of LocAgent with dense retrieval methods + LLM reranker for Python repositories in SWE-PolyBench
 Dataset Model Recall@3 Accuracy@3 MRR@3
-SWEBench-Verified LocAgent 0.18 0.17 0.12
+SWE-PolyBench LocAgent 0.18 0.17 0.12
 SweRankEmbed-Small ZS+ SpIDER 0.44 0.36 0.54
-+ Reranker
++ LLM reranker
 CodeSAGE-Small ZS + SpIDER 0.23 0.18 0.31
-+ Reranker
++ LLM reranker
 CodeSAGE-Small-PEFT + SpIDER 0.32 0.25 0.42
-+ Reranker
++ LLM reranker
 BM25+ SpISR 0.30 0.23 0.38
-SWE-PolyBench LocAgent 0.23 0.17 0.21
-SweRankEmbed-Small ZS+ SpIDER 0.61 0.56 0.63
-+ Reranker
-CodeSAGE-Small ZS + SpIDER 0.– 0.– 0.–
-+ Reranker
-CodeSAGE-Small-PEFT + SpIDER 0.52 0.46 0.56
-+ Reranker
-BM25+ SpISR 0.– 0.– 0.–
-+ Reranker
 Table 10: SWE-PolyBench retrieval performance with 95% confidence intervals ($\mu\pm$ CI) for $K=20$, $N=500$, $C=5$, $d=4$
@@
 SR 0.07 $\pm$ 0.03 0.05 $\pm$ 0.03
-Table 12: SpIDER + SweRankEmbed-Small performance with varying neighbor filtering threshold NN for $K=20$, $C=3$, $d=4$.
+Table 11: Ablation on primary neighborhood filtering threshold NN using SpIDER + SweRankEmbed-Small for $K=20$, $C=3$, $d=4$.
 Increasing NN improves coverage and Recall@20 with modest increases in LLM token usage.
@@
 Bottom row: Accuracy.
+Table 14: Comparison of dense retrieval and dense retrieval + LLM reranker for $K=20$ and $N=500$, $C=5$, $d=4$ on SWE-PolyBench. Winners per metric per embedding model are highlighted in light blue.
+Language Model Retrieval Dense Retrieval Only Dense + LLM reranker
+Recall@20 Accuracy@20 MRR@20 Recall@3 Accuracy@3 MRR@3
+Python SweRankEmbed-Small ZS SpIDER 0.49 0.40 0.30 0.44 0.36 0.54
+DER 0.42 0.35 0.26 0.38 0.31 0.48
+CodeSAGE-Small ZS SpIDER 0.27 0.21 0.13 0.23 0.18 0.31
+DER 0.21 0.17 0.11 0.19 0.16 0.25
+CodeSAGE-Small-PEFT SpIDER 0.37 0.27 0.24 0.32 0.25 0.42
+DER 0.29 0.20 0.21 0.26 0.20 0.38
+BM25 SpISR 0.33 0.26 0.19 0.30 0.23 0.38
+SR 0.31 0.24 0.18 0.27 0.22 0.35
+Java SweRankEmbed-Small ZS SpIDER 0.36 0.27 0.18 0.25 0.17 0.36
+DER 0.31 0.24 0.18 0.24 0.16 0.33
+CodeSAGE-Small ZS SpIDER 0.23 0.17 0.10 0.16 0.09 0.26
+DER 0.15 0.10 0.07 0.11 0.07 0.19
+CodeSAGE-Small-PEFT SpIDER 0.28 0.18 0.15 0.20 0.14 0.28
+DER 0.21 0.14 0.13 0.17 0.11 0.25
+BM25 SpISR 0.16 0.10 0.10 0.12 0.08 0.20
+SR 0.15 0.10 0.10 0.12 0.09 0.19
+JavaScript SweRankEmbed-Small ZS SpIDER 0.60 0.52 0.18 0.43 0.37 0.43
+DER 0.52 0.45 0.16 0.41 0.34 0.42
+CodeSAGE-Small ZS SpIDER 0.38 0.32 0.09 0.30 0.25 0.31
+DER 0.26 0.21 0.07 0.22 0.17 0.24
+CodeSAGE-Small-PEFT SpIDER 0.55 0.48 0.15 0.41 0.35 0.41
+DER 0.45 0.38 0.14 0.36 0.30 0.39
+BM25 SpISR 0.46 0.39 0.13 0.36 0.30 0.38
+SR 0.36 0.29 0.11 0.31 0.25 0.34
+TypeScript SweRankEmbed-Small ZS SpIDER 0.35 0.28 0.06 0.27 0.21 0.29
+DER 0.29 0.21 0.18 0.25 0.19 0.30
+CodeSAGE-Small ZS SpIDER 0.16 0.10 0.03 0.14 0.09 0.19
+DER 0.10 0.06 0.06 0.09 0.06 0.14
+CodeSAGE-Small-PEFT SpIDER 0.25 0.18 0.04 0.22 0.16 0.27
+DER 0.19 0.13 0.11 0.18 0.12 0.24
+BM25 SpISR 0.11 0.09 0.02 0.10 0.07 0.11
+SR 0.07 0.05 0.04 0.07 0.05 0.09
+Table 15: Dense retrieval performance across benchmarks for $K=20$, $N=500$, $C=5$, $d=4$. Winners per metric per embedding model highlighted in light blue.
+Language Model Retrieval Recall@20 Accuracy@20 MRR@20
+Multi-SWEBench
+Java SweRankEmbed-Small ZS SpIDER 0.37 0.28 0.08
+DER 0.32 0.24 0.08
+CodeSAGE-Small ZS SpIDER 0.21 0.19 0.05
+DER 0.15 0.13 0.03
+CodeSAGE-Small-PEFT SpIDER 0.31 0.25 0.10
+DER 0.23 0.17 0.09
+BM25 SpISR 0.23 0.17 0.07
+SR 0.22 0.17 0.05
+JavaScript SweRankEmbed-Small ZS SpIDER 0.39 0.30 0.18
+DER 0.31 0.23 0.15
+CodeSAGE-Small ZS SpIDER 0.17 0.12 0.08
+DER 0.09 0.06 0.05
+CodeSAGE-Small-PEFT SpIDER 0.32 0.24 0.13
+DER 0.25 0.18 0.11
+BM25 SpISR 0.21 0.14 0.08
+SR 0.14 0.10 0.06
+TypeScript SweRankEmbed-Small ZS SpIDER 0.52 0.42 0.08
+DER 0.40 0.32 0.23
+CodeSAGE-Small ZS SpIDER 0.33 0.26 0.04
+DER 0.26 0.21 0.13
+CodeSAGE-Small-PEFT SpIDER 0.30 0.26 0.05
+DER 0.19 0.15 0.12
+BM25 SpISR 0.32 0.27 0.04
+SR 0.24 0.21 0.11
+SWEBench-Verified
+Python SweRankEmbed-Small ZS SpIDER 0.61 0.56 0.32
+DER 0.54 0.49 0.29
+CodeSAGE-Small ZS SpIDER 0.30 0.27 0.13
+DER 0.21 0.19 0.10
+CodeSAGE-Small-PEFT SpIDER 0.55 0.49 0.27
+DER 0.43 0.38 0.22
+BM25 SpISR 0.45 0.41 0.25
+SR 0.38 0.34 0.22
+Table 16: Impact of the number of center nodes ($C$) on retrieval performance using SpIDER + SweRankEmbed-Small.
+Increasing $C$ generally improves Recall@20 and Acc@20 across languages, while MRR@20 shows modest variation.
+$\boldsymbol{C}$ Python JavaScript Java TypeScript
+Recall@20 Acc@20 MRR@20 Recall@20 Acc@20 MRR@20 Recall@20 Acc@20 MRR@20 Recall@20 Acc@20 MRR@20
+1 0.47 0.38 0.29 0.56 0.48 0.18 0.32 0.24 0.18 0.32 0.24 0.06
+3 0.49 0.39 0.30 0.59 0.51 0.18 0.35 0.26 0.19 0.31 0.24 0.06
+5 0.49 0.40 0.30 0.60 0.52 0.18 0.36 0.27 0.18 0.35 0.28 0.06
+7 0.49 0.39 0.30 0.60 0.52 0.18 0.36 0.27 0.19 0.37 0.30 0.06
+9 0.48 0.39 0.30 0.60 0.53 0.19 0.37 0.27 0.19 0.39 0.31 0.06
+MRR@$K$ Gains Depend on Graph Neighborhoods
+The MRR@$K$ metric benefits from SpIDER primarily when the first relevant ground-truth item within the top-$K$ results is a neighbor of a center node.
+In cases where the first relevant item is a non-center node, particularly in TypeScript as shown in Tab.
+2SpIDER may rank neighbors of center nodes ahead, slightly reducing MRR@$K$.
+Since SpIDER is not explicitly designed to optimize the precise ranking order within the top-$K$ results, improvements in MRR@$K$ over baseline DER methods are modest but still observable.